### v1

# Stochastic Gradient Descent Jittering for Inverse Problems: Alleviating the Accuracy-Robustness Tradeoff

Abstract
Inverse problems aim to reconstruct unseen data from corrupted or perturbed measurements. While most work focuses on improving reconstruction quality, generalization accuracy and robustness are equally important, especially for safety-critical applications. Model-based architectures (MBAs), such as loop unrolling methods, are considered more interpretable and achieve better reconstructions. Empirical evidence suggests that MBAs are more robust to perturbations than black-box solvers, but the accuracy-robustness tradeoff in MBAs remains underexplored. In this work, we propose a simple yet effective training scheme for MBAs, called SGD jittering, which injects noise iteration-wise during reconstruction. We theoretically demonstrate that SGD jittering not only generalizes better than the standard mean squared error training but is also more robust to average-case attacks. We validate SGD jittering using denoising toy examples, seismic deconvolution, and single-coil MRI reconstruction. The proposed method achieves cleaner reconstructions for out-of-distribution data and demonstrates enhanced robustness to adversarial attacks.
Introduction
Inverse problems (IPs) aim to recover the underlying signal ${\bm{x}}\in\mathbb{R}^{n}$ from corrupted observation ${\bm{y}}\in\mathbb{R}^{m}$, where $m$ is often less than $n$, leading to non-unique solutions. IPs are typically modeled as:
${\bm{y}}=\bm{A}{\bm{x}}+{\bm{z}},$ (1)
where $\bm{A}\in\mathbb{R}^{m\times n}$ is a known forward model for linear IPs, and ${\bm{z}}$ represents unknown noise in the observation.
Machine learning (ML) algorithms show significantly better reconstructions than the classical optimization methods. The majority of the work in the field aims to promote reconstruction quality using different deep-learning models, such as transformers (Fabian, Tinaz, and Soltanolkotabi 2022) and diffusion models (Whang et al. 2022). In particular, model-based architectures (MBAs) such as loop unrolling (LU) networks inspired by classical optimization methods, unroll the objective function into a sequence of trainable iterative steps with the forward model $\bm{A}$ involved in every iteration. MBAs are considered to be more interpretable than neural networks that learn a direct inverse without using $\bm{A}$ (Monga, Li, and Eldar 2021), and achieve state-of-the-art reconstructions in different fields, such as natural image restoration (Gilton, Ongie, and Willett 2021) and medical image reconstruction (Bian 2024).
While reconstruction quality is crucial, robustness and generalization accuracy are equally important for IPs. Robustness refers to the model’s ability to handle unwanted noise, while generalization accuracy measures how well the model performs on unseen data. For example, in medical image reconstruction, a model must generalize well to detect subtle details like small bone fractures and be robust enough to eliminate artifacts caused by measurement noise. Despite their importance, robustness and generalization have not been thoroughly explored in the context of IPs, particularly in MBAs. While some studies provide empirical assessments of robustness (Gilton, Ongie, and Willett 2021; Genzel, Macdonald, and März 2022), there is a noticeable lack of theoretical analysis (Gottschling et al. 2020).
In this work, we model MBA training as a bilevel optimization problem without making specific assumptions about neural network architectures. We theoretically analyze its generalization and robustness risks in solving denoising problems and demonstrate that standard mean-squared error (MSE) training may perform poorly in both metrics.
Technologies like adversarial training (AT) (Fawzi, Moosavi-Dezfooli, and Frossard 2016; Ford et al. 2019; Kannan, Kurakin, and Goodfellow 2018) and input noise injection (Fawzi, Moosavi-Dezfooli, and Frossard 2016; Ford et al. 2019; Kannan, Kurakin, and Goodfellow 2018) are commonly used to enhance robustness, while approaches such as weight perturbation (Wu, Xia, and Wang 2020), stochastic training (Hardt, Recht, and Singer 2016) and sharpness-aware minimization (Foret et al. 2020) are designed to improve generalization. However, there is often a tradeoff between robustness and accuracy (Tsipras et al. 2018; Zhang et al. 2023; Gottschling et al. 2020; Krainovic, Soltanolkotabi, and Heckel 2023). Although strategies like careful loss function design (Pang et al. 2022; Kannan, Kurakin, and Goodfellow 2018; Zhang et al. 2019; Zheng, Zhang, and Mao 2021), data augmentation (Zhang et al. 2017), and robust architectural choices such as dropout (Yang et al. 2020) and teacher-student models (Zhao, Wang, and Wei 2024) have been explored to mitigate this tradeoff in classification tasks, the challenge of balancing accuracy and robustness in IP solvers still remains an open question (Gottschling et al. 2020). Recent work by Krainovic, Soltanolkotabi, and Heckel (2023) introduces jittering in ${\bm{y}}$ for black-box IP neural networks, achieving a comparable level of worst-case robustness to AT but with significant accuracy loss.
To bridge this gap, this work aims to improve robustness through a novel training scheme for MBAs, while maintaining generalization accuracy for IPs. Note that while specially designed architectures can also improve robustness or generalization, they are not discussed here as this work focuses on investigating different training schemes.
Our Contributions
In this work, we propose a stochastic gradient descent (SGD) jittering scheme in training MBAs. This method unrolls the objective function for solving IPs using a gradient descent algorithm, with small, random noises added to the gradient updates at each iteration to improve the training process. During inference, these jittering noises are removed, allowing the model to perform as it would in standard evaluation settings. This technique has demonstrated greater robustness and better generalization than traditional MSE training, while also being more time-efficient than AT. Our contribution can be summarized as follows,
  • (network assumptions) We investigate the performance of MBAs with general non-convex neural networks, where training often results in non-unique parameters. This analysis is crucial for demonstrating improvements across various metrics, particularly in robustness and generalization accuracy.
  • (robustness) We theoretically prove that SGD jittering implicitly promotes average-case robustness than regular MSE training. In the experiments, we further demonstrate the effectiveness of SGD jittering under worst-case attacks.
  • (generalization) We are the first to analyze generalization accuracy for IPs with small perturbations in the testing data, rather than relying on PAC-Bayes Theorem, which hypothetically promotes generalization through Hessian regularization.
  • (convergence) We also demonstrate that under the assumption of Lipschitz continuity, SGD jittering converges, ensuring the training is feasible.
Related Work
Analysis of Robustness
Smoothness is closely related to the robustness of neural networks. Works like Salman et al. (2019); Cohen, Rosenfeld, and Kolter (2019); Lecuyer et al. (2019); Li et al. (2018) show that regularization in smoothness promotes the robustness of classifiers. On the other hand, the analysis of IP is very different from classifications. Krainovic, Soltanolkotabi, and Heckel (2023) analyzes the worst-case robustness for IPs using a simple linear reconstruction model and directly solves for robustness risk. However, optimizing a convex model results in a solution that excels in only one metric (accuracy or robustness). In this work, we examine nonlinear MBAs in general forms to show improvements in both metrics.
Analysis of Generalization
Demonstrating better generalization of classifiers often involves showing larger classification margins (Cao et al. 2019; Kawaguchi, Kaelbling, and Bengio 2017; Lim et al. 2021), but the proof for IPs is very different. A common but indirect approach shows an algorithm can regularize the Hessian of loss landscape with respect to the network parameters. Models that find flat minima in the loss landscape tend to generalize better in practice (Jiang et al. 2019; Keskar et al. 2016; Neyshabur et al. 2017; Foret et al. 2020). Many works demonstrate regularization of the Hessian (Xie, Sato, and Sugiyama 2020; Keskar et al. 2016; Foret et al. 2020) then complete the proof using the PAC-Bayes Theorem in hypothesis. While the PAC-Bayes theorem provides a bound on the generalization error, it does not directly incorporate curvature measures. The exact theoretical mechanism behind it is still not fully understood. In this work, instead of connecting to the PAC-Bayes Theorem, we rewrite the generalization risk and demonstrate that SGD jittering results in a smaller generalization risk than regular MSE training.
Noise Injections
Noise injection can be viewed as implicit regularization during training, which is employed to enhance model robustness and generalization. Input injection adds random noises to the signal space, which can be viewed as a data augmentation strategy. It is widely studied in classifications (Fawzi, Moosavi-Dezfooli, and Frossard 2016; Ford et al. 2019; Rusak et al. 2020), and recently shown to promote robustness in IPs (Krainovic, Soltanolkotabi, and Heckel 2023). Weight injection instead perturbed network parameters during training, which is shown to help escape from local minima (Zhu et al. 2018; Nguyen et al. 2019) and promote smooth solutions (Orvieto et al. 2023; Liu et al. 2021; Orvieto et al. 2022). Layer-wise noise injection (Lim et al. 2021) in a recurrent network has been shown to generalize better for classifiers. Weights can also be perturbed before parameter updates to enhance generalization, as in stochastic gradient descent (SGD). We will explore SGD in detail in the next subsection.
Connection to SGD Training
SGD training is effective in promoting generalization through implicit regularization (Smith et al. 2021). In large overparameterized models with non-unique global solutions, SGD favors networks with flat minima (Hochreiter and Schmidhuber 1997). The implicit regularization properties of SGD have been explored in contexts such as linear regression (Zou et al. 2021), logistic regression (Ji and Telgarsky 2018), and convolutional networks (Gunasekar et al. 2018). Unlike SGD training, which adds noise to network parameters, the proposed SGD jittering method introduces noise at each iterative update of the reconstructions in the lower-level optimization. Despite its name, SGD jittering is structurally similar to layer-wise noise injection within MBAs.
Definition of Robustness and Generalization
We formally define average-case robustness and generalization risks for a learned inverse mapping $H_{\theta}:{\bm{y}}\mapsto{\bm{x}}$ in IPs. Average-case robustness is also known as the “natural” robustness for average performance (Rice et al. 2021; Hendrycks and Dietterich 2019), which is equally important as in worst-case. During evaluation, a small vector ${\bm{g}}$ sampled from the distribution $\mathcal{P}_{g}$ is added to a data point ${\bm{x}}\in\mathcal{D}$. The new ground-truth becomes ${\bm{x}}_{g}={\bm{x}}+{\bm{g}}$, with the corresponding observation ${\bm{y}}_{g}=\bm{A}{\bm{x}}_{g}+{\bm{z}}={\bm{y}}+\bm{A}{\bm{g}}$.
Definition 1.
The average-case robustness risk of $H_{\theta}$ measures the distance between the unperturbed ${\bm{x}}$ to the recovered signal $\hat{{\bm{x}}}_{g}=H_{\theta}({\bm{y}}_{g})$ with the presence of ${\bm{g}}$,
$R_{e}(\theta)=\mathbb{E}_{{\bm{x}},{\bm{y}}\in\mathcal{D},{\bm{g}}\in\mathcal{P}_{g}}||{\bm{x}}-H_{\theta}({\bm{y}}_{g})||^{2}_{2}.$ (2)
Definition 2.
The generalization risk of $H_{\theta}$ measures the distance between the actual ground-truth ${\bm{x}}_{g}$ to the recovered signal $\hat{{\bm{x}}}_{g}=H_{\theta}({\bm{y}}_{g})$ with the presence of ${\bm{g}}$,
$\mathcal{G}(\theta)=\mathbb{E}_{{\bm{x}},{\bm{y}}\in\mathcal{D},{\bm{g}}\in\mathcal{P}_{g}}||{\bm{x}}_{g}-H_{\theta}({\bm{y}}_{g})||^{2}_{2}.$ (3)
Robustness measures how close the reconstruction is compared to the data within the dataset, and the generalization accuracy measures how well the reconstruction obeys the physics of the inverse problem even for out-of-distribution (OOD) data. Defining generalization in regression is not straightforward, but for IPs, we can achieve this using the underlying forward models. Typically, a tradeoff between robustness and accuracy exists for classical feedforward neural networks (Zhang et al. 2019; Yang et al. 2020; Stutz, Hein, and Schiele 2019). We aim to illustrate that the SGD jittering technique enhances robustness and accuracy for MBAs without incorporating explicit regularization.
Model-based Architecture Overview
MBAs are inspired by classical optimizations, where the underlying signal ${\bm{x}}$ can be estimated via,
$\hat{{\bm{x}}}=\operatorname*{arg\,min}_{\bm{x}}\frac{1}{2}||{\bm{y}}-\bm{A}{\bm{x}}||_{2}^{2}+r({\bm{x}}),$
where $r$ is an arbitrary regularization function to ensure the uniqueness of the reconstruction. When $r$ is differentiable, the optimal solution can be solved via gradient descent (GD), iterating through steps $k=1,2,3...$ with a constant step-size $\eta$,
${\bm{x}}_{k+1}={\bm{x}}_{k}-\eta\bm{A}^{\top}(\bm{A}{\bm{x}}_{k}-{\bm{y}})+\eta\nabla r({\bm{x}}_{k}).$
In a loop unrolling (LU) architecture using GD, $\nabla r$ is replaced by a shared-weight neural network, denoted by $f_{\theta}$. The process begins with an initial estimate ${\bm{x}}_{0}$, which is iteratively updated with a finite number of iterations. The final output ${\bm{x}}_{K}$ represents the predicted reconstruction and is compared to the desired target. Different variants of LU are summarized in Monga, Li, and Eldar (2021).
MBA Training Scheme Overview
In this section, we summarize some common training schemes for general regression problems, adapted here for MBA training. Following this, we introduce the proposed SGD jittering method for training MBAs.
Regular MSE Training
MSE training is the most commonly used training scheme for training regression problems and MBAs, where the goal using MBAs is to minimize the following risk w.r.t. $\theta$,
$R(\theta)$ $=\mathbb{E}_{({\bm{x}},{\bm{y}})\in\mathcal{D}}||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\quad\textrm{ where, }$, (4)
$\hat{{\bm{x}}}$ $=\operatorname*{arg\,min}_{x}~{}\frac{1}{2}\|{\bm{y}}-\bm{A}{\bm{x}}\|^{2}_{2}+r_{\theta}({\bm{x}})$
Notice that MBAs do to learn $r$ directly, but instead learn $\nabla r$ using a neural network $f_{\theta}$. We denote $r_{\theta}$ to highlight its dependency of $\theta$. The goal of MSE training is to minimize the squared distance between the output ${\bm{x}}_{K}$ and the ground-truth ${\bm{x}}$. MSE training might overfit the patterns in training data, making it more vulnerable to variations in ${\bm{y}}$.
Adversarial Training
AT is widely recognized as one of the most effective methods for training robust neural networks (Wang et al. 2021). AT finds the worst-case attack for each training instance and adds it to the input data designed to deceive the model. By learning from these challenging examples, AT improves the model’s robustness. AT aims to minimize the worst-case robustness risk for MBAs as follows,
$R_{\epsilon}(\theta)=\mathbb{E}_{({\bm{x}},{\bm{y}})\in\mathcal{D}}\left[\begin{aligned} &\max_{{\bm{e}}:||{\bm{e}}||_{2}\leq\epsilon}\quad||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\\ &\textrm{s.t. }\hat{{\bm{x}}}=\operatorname*{arg\,min}_{x}~{}\frac{1}{2}\|{\bm{y}}+{\bm{e}}-\bm{A}{\bm{x}}\|^{2}_{2}+r_{\theta}({\bm{x}})\end{aligned}\right].$ (5)
Fast gradient sign method (FGSM) and projected gradient descent (PGD) are methods to find the attack vector ${\bm{e}}$, where (Athalye, Carlini, and Wagner 2018) demonstrate that training a defense model with PGD adversarial attacks is more robust. However, AT is slow as it requires iterative solvers for the attack vector (Shafahi et al. 2019).
AT might suffer from bad generalizations. For example, in noiseless IPs, ${\bm{y}}=\bm{A}{\bm{x}}$, the true inverse $\bm{A}^{-1}$ exists. AT learns an inverse mapping $H_{\theta}$ so that $H_{\theta}({\bm{y}}+{\bm{e}})={\bm{x}}$, but the true estimated should be $\bm{A}^{-1}({\bm{y}}+{\bm{e}})\neq{\bm{x}}$.
Input Jittering
Noise injection to the input is widely used in general ML tasks to promote robustness (Salman et al. 2019; Cohen, Rosenfeld, and Kolter 2019; Lecuyer et al. 2019). In IPs that aim to map from ${\bm{y}}$ to ${\bm{x}}$, (Krainovic, Soltanolkotabi, and Heckel 2023) proposes to add zero-mean random Gaussian vector ${\bm{w}}$ in ${\bm{y}}$ to a neural network that learns a direct reconstruction, and demonstrate that with careful selection of the jittering variance, this training scheme can attain a comparable level of worst-case robustness to that achieved by AT. In MBAs, the input jittering approach is expressed as,
$J^{I}_{\sigma_{w}}(\theta)$ $=\mathbb{E}_{{\bm{w}},({\bm{x}},{\bm{y}})\in\mathcal{D}}||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\quad\textrm{ where,}$, (6)
$\hat{{\bm{x}}}$ $=\operatorname*{arg\,min}_{x}~{}\frac{1}{2}\|{\bm{y}}+{\bm{w}}-\bm{A}{\bm{x}}\|^{2}_{2}+r_{\theta}({\bm{x}}).$
This method learns an inverse mapping from ${\bm{y}}+{\bm{w}}$ to ${\bm{x}}$, which might suffer from generalization issues in the same way as AT. In fact, Krainovic, Soltanolkotabi, and Heckel (2023) demonstrates a significant accuracy loss using input jittering for black-box neural networks.
Proposed Training Scheme
Another line of work injects random noises layer-wise into the neural network, where (Hodgkinson et al. 2021; Jim, Giles, and Horne 1996; Liu et al. 2020) model the layer-wise injection as stochastic differential equation and show implicit regularization in robustness, and (Lim et al. 2021) added noises to the recurrent neural network and demonstrates better generalization for classification tasks.
In this work, we propose adding random noise vectors ${\bm{w}}_{k}$ drawn from a zero-mean Gaussian distribution to the learned gradient updates $f_{\theta}({\bm{x}}_{k})$, with the noise being re-sampled independently for each iteration $k=1,2,3,...$. The noisy gradients are unbiased, $\mathbb{E}[f_{\theta}({\bm{x}})+{\bm{w}}_{k}]=\mathbb{E}[f_{\theta}({\bm{x}})]$ for all ${\bm{x}}$. Let $\bar{{\bm{w}}}=\{{\bm{w}}_{1},...,{\bm{w}}_{K}\}$. The estimated reconstruction is solved iteratively using this noisy gradient descent which is essentially the stochastic gradient descent (SGD) with step size $\eta$. Thus, the learning objective is as follows,
$J^{SGD}_{\sigma_{w_{k}}}(\theta)$ $=\mathbb{E}_{\bar{{\bm{w}}},({\bm{x}},{\bm{y}})\in\mathcal{D}}||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\quad\textrm{ where,}$, (7)
$\hat{{\bm{x}}}$ $=\lim_{k\rightarrow\infty}{\bm{x}}_{k}-\eta\big{(}\bm{A}^{\top}(\bm{A}{\bm{x}}_{k}-{\bm{y}})+f_{\theta}({\bm{x}}_{k})+{\bm{w}}_{k}\big{)}$
In AT and input jittering training, the noise vectors are added to ${\bm{y}}$ throughout the entire inversion, introducing a bias to the reconstruction compared to the actual ground-truth. In contrast, SGD jittering adds noise to the learned gradient descent updates. The lower-level problem in SGD jittering minimizes $\frac{1}{2}||{\bm{y}}-\bm{A}{\bm{x}}||_{2}^{2}+r_{\theta}({\bm{x}})$, with the convergence result in Lemma 1 ensuring the correct inverse mapping even with noise in gradient descent.
We modified the convergence result of the SGD in (Ghadimi and Lan 2013) for lower-level SGD in MBAs as follows, where the proof of Lemma 1 can be found in the Appendix.
Assumption 1.
(SGD noises) Assume noises ${\bm{w}}_{k}\sim\mathcal{N}(0,\sigma_{w_{k}}^{2}/n\mathbf{I})$ sampled from i.i.d. zero-mean Gaussian distribution for each $k=1,2,3,...$ are added to each gradient descent iteration, so that $\mathbb{E}||{\bm{w}}_{k}||_{2}^{2}=\sigma_{w_{k}}^{2}$.
Lemma 1.
(Convergence of MBAs-SGD for IPs, modified from SGD convergence result in (Ghadimi and Lan 2013)) Let $F({\bm{x}})=\frac{1}{2}||{\bm{y}}-\bm{A}{\bm{x}}||_{2}^{2}+r({\bm{x}})$ denote the lower-level objective function. Assume $f_{\theta}({\bm{x}})$ is $L$-Lipschitz continuous,
$||f_{\theta}({\bm{x}}+\Delta)-f_{\theta}({\bm{x}})||_{2}\leq L||\Delta||_{2},\quad\forall{\bm{x}},\Delta$
With assumption 1, and step-sizes $\eta_{k}<2/L$, SGD in minimizing $F({\bm{x}})$ converges to a stationary point.
SGD noise also implicitly enhances robustness. Intuitively, exposing a neural network to noisier inputs during training encourages it to converge to smoother regions of the loss landscape, which are less sensitive to input changes. This leads to more stable outputs that are less affected by perturbations during evaluation. We will analyze how SGD jittering contributes to this robustness in the following section.
Analysis of Denoising Problem
In this section, we analyze the generalization and robustness of SGD jittering in addressing the denoising problem. We begin by outlining the relevant assumptions.
Assumption 2.
(Denoising) The forward model in denoising is identity mapping, so that ${\bm{y}}={\bm{x}}+{\bm{z}}$. Signals ${\bm{x}}\in\mathbb{R}^{n}$ are sampled from distribution $\mathcal{X}$, and noises are sampled i.i.d. from ${\bm{z}}\sim\mathcal{N}(0,\sigma_{z}^{2}/n\bm{I})$.
Assumption 3.
(Neural Network) Assume $f_{\theta}$ is twice differentiable with respect to the input, so the Hessian exists.
Figure 1: 2D point denoising example using MSE training, AT, and SGD jittering. Each method is evaluated using regular in-distribution data, under average- and worst-case attacks to ${\bm{y}}$ and with additional bias (from left to right columns respectively).
Main Theoretical Results
Given the assumptions, we state the first theorem regarding generalization accuracy. Let $H_{\theta}$ denote the inverse process learned using MBAs for the rest of the analysis.
Theorem 2.
Assuming a small zero-meam Gaussian random vector ${\bm{g}}$ with $\mathbb{E}||{\bm{g}}||^{2}_{2}=\sigma_{g}^{2}$ is added to data ${\bm{x}}\in\mathcal{D}$ during evaluation. The generalization risk for denoising problem is defined as,
$\mathcal{G}(\theta)=\mathbb{E}_{{\bm{x}},{\bm{y}}\in\mathcal{D},{\bm{g}}}||{\bm{x}}+{\bm{g}}-H_{\theta}({\bm{y}}+{\bm{g}})||^{2}_{2}.$ (8)
Under assumptions 1, 2 and 3, SGD jittering generalizes better than regular MSE training,
$\mathcal{G}(\theta^{sgd})\leq\mathcal{G}(\theta^{mse}).$
Proof.
For any trained model $\theta$, we have generalization risk in (8). We can express $H_{\theta}({\bm{y}}+{\bm{g}})$ in terms of $H_{\theta}({\bm{y}})$ for the same set of parameters via iterative expansions. For iterations $k=0,1,...,K$, let the sequence $\{{\bm{x}}_{0},{\bm{x}}_{1},...,{\bm{x}}_{K}\}$ denote the denoising trajectory from observation ${\bm{y}}+{\bm{g}}$, where $H_{\theta}({\bm{y}}+{\bm{g}})={\bm{x}}_{K}$, and let $\{{\bm{x}}_{0}^{\prime},{\bm{x}}_{1}^{\prime},...,{\bm{x}}_{K}^{\prime}\}$ denote the denoising trajectory from observation ${\bm{y}}$ where $H_{\theta}({\bm{y}})={\bm{x}}_{K}^{\prime}$. Then we can rewrite the reconstruction trajectory with the appearance of ${\bm{g}}$ as follows,
${\bm{x}}_{0}={\bm{x}}_{0}^{\prime}+{\bm{g}}={\bm{y}}+{\bm{g}}$
${\bm{x}}_{1}=(1-\eta){\bm{x}}_{0}+\eta({\bm{y}}+{\bm{g}})-\eta f_{\theta}({\bm{x}}_{0})$
$=\big{(}(1-\eta){\bm{x}}_{0}^{\prime}+\eta{\bm{y}}-\eta f_{\theta}({\bm{x}}_{0}^{\prime})\big{)}+\big{(}{\bm{g}}+\eta f_{\theta}({\bm{x}}_{0}^{\prime})-\eta f_{\theta}({\bm{x}}_{0})\big{)}$
$={\bm{x}}_{1}^{\prime}+{\bm{g}}+\eta(f_{\theta}({\bm{x}}_{0}^{\prime})-f_{\theta}({\bm{x}}_{0}))$
${\bm{x}}_{2}=(1-\eta){\bm{x}}_{1}+\eta({\bm{y}}+{\bm{g}})-\eta f_{\theta}({\bm{x}}_{1})$
$={\bm{x}}_{2}^{\prime}+{\bm{g}}+\eta(1-\eta)(f_{\theta}({\bm{x}}_{0}^{\prime})-f_{\theta}({\bm{x}}_{0}))+\eta(f_{\theta}({\bm{x}}_{1}^{\prime})-f_{\theta}({\bm{x}}_{1}))$
$\quad$...
${\bm{x}}_{k+1}={\bm{x}}_{k+1}^{\prime}+{\bm{g}}+\sum_{i=0}^{k}\eta(1-\eta)^{k-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}))$
Let $K=k+1$ and plug in the above expansion to (8). For brevity, we omit the subscripts of expectations in the proof.
$\mathcal{G}(\theta)=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}-\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}))||^{2}_{2}$
$=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}))||^{2}_{2}$
The first term is the testing accuracy when ${\bm{x}},{\bm{y}}$ are in the data distribution, while the second term is affected by ${\bm{g}}$. MSE training minimizes the first term solely, but no extra regularization on the second term. On the other hand, we will show that SGD jittering training also penalizes the magnitude of the second term, thus obtaining a smaller generalization risk than MSE training.
We view the SGD jittering process as a noisy version of regular MSE training, the goal is to write the SGD jittering risk in terms of the noiseless updates with some implicit regularization. Let $\{{\bm{x}}_{0}^{\prime},{\bm{x}}_{1}^{\prime},...,{\bm{x}}_{K}^{\prime}\}$ denote the noiseless GD trajectory of a model $\theta$. We write the iterative updates of SGD jittering training $\{{\bm{x}}_{0}^{sgd},{\bm{x}}_{1}^{sgd},...,{\bm{x}}_{K}^{sgd}\}$ for the same set of parameters $\theta$ in terms of ${\bm{x}}_{k}^{\prime}$s, which tells the deviation of each intermediate reconstruction between the SGD jittering and its noiseless counterpart,
${\bm{x}}_{0}^{sgd}={\bm{y}}={\bm{x}}_{0}^{\prime}+{\bm{w}}_{0}$
${\bm{x}}_{1}^{sgd}=(1-\eta){\bm{x}}_{0}^{sgd}+\eta{\bm{y}}-\eta f_{\theta}({\bm{x}}_{0}^{sgd})-\eta{\bm{w}}_{1}$
$={\bm{x}}_{1}^{\prime}+\eta(f_{\theta}({\bm{x}}_{0}^{\prime})-f_{\theta}({\bm{x}}_{0}^{sgd}))-(\eta(1-\eta){\bm{w}}_{0}+\eta{\bm{w}}_{1})$
${\bm{x}}_{2}^{sgd}=(1-\eta){\bm{x}}_{1}^{sgd}+\eta{\bm{y}}-\eta f_{\theta}({\bm{x}}_{1}^{sgd})-\eta{\bm{w}}_{2}$
$={\bm{x}}_{2}^{\prime}+\eta\big{(}f_{\theta}({\bm{x}}_{1}^{\prime})-f_{\theta}({\bm{x}}_{1}^{sgd})\big{)}+\eta(1-\eta)\big{(}f_{\theta}({\bm{x}}_{0}^{\prime})-f_{\theta}({\bm{x}}_{0}^{sgd})\big{)}$
$\qquad-\big{(}\eta(1-\eta)^{2}{\bm{w}}_{0}+\eta(1-\eta){\bm{w}}_{1}+\eta{\bm{w}}_{2}\big{)}$
${\bm{x}}_{k+1}^{sgd}=(1-\eta){\bm{x}}_{k}^{sgd}+\eta{\bm{y}}-\eta f_{\theta}({\bm{x}}_{k}^{sgd})-\eta{\bm{w}}_{k+1}$
$={\bm{x}}_{k+1}^{\prime}+\sum_{i=0}^{k}\eta(1-\eta)^{k-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd}))$
$\qquad-\sum_{i=0}^{k+1}\eta(1-\eta)^{k+1-i}{\bm{w}}_{i}.$
Let $K=k+1$, SGD jittering training risk is,
$\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{sgd}||_{2}^{2}=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}~{}||\sum_{i=0}^{K}\eta(1-\eta)^{K-i}{\bm{w}}_{i}||_{2}^{2}$
$+\mathbb{E}~{}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd}))||_{2}^{2}$
$-2\mathbb{E}\left[\langle\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}f_{\theta}({\bm{x}}_{i}^{sgd}),\sum_{i=0}^{K}\eta(1-\eta)^{K-i}{\bm{w}}_{i}\rangle\right].$
The first component corresponds to the MSE loss when the trajectory is not perturbed by jittering noise. The second term is weighted noise variance independent of the network parameters. The third term adds extra regularization to the difference in outputs of $f_{\theta}$ between noisy and clean trajectories. It penalizes more as the iterations approach the final output. The last term is the cross product between the jittering noise and the perturbed reconstructions from $f$, which is lower bounded since the overall SGD jittering risk is non-negative.
When choosing $\sigma_{{\bm{w}}_{k}}^{2}$ at iteration $k$ such that $\sum_{i=0}^{k}\eta^{2}(1-\eta)^{2(k-i)}\sigma_{w_{i}}^{2}=\sigma_{\bm{g}}^{2}$, the reconstruction from ${\bm{x}}+{\bm{g}}$ is a special case in SGD jittering training where ${\bm{w}}_{k}$ are independently sampled at each iteration. Thus, minimizing the regularization term $\mathbb{E}_{({\bm{x}},{\bm{y}})\sim\mathcal{D},{\bm{w}}_{k}\forall k}~{}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd}))||_{2}^{2}$ also reduces the penalty term in computing generalization accuracy $\mathbb{E}_{{\bm{x}},{\bm{y}}\in\mathcal{D},{\bm{g}}}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}))||^{2}_{2}$. Therefore, $\mathcal{G}(\theta^{sgd})\leq\mathcal{G}(\theta^{mse})$ due to the implicit regularization term in SGD jittering training.
In addition, we can expand the regularization term in SGD training via Taylor expansion. Let $\bm{\delta}_{k+1}=\sum_{i=0}^{k}\eta(1-\eta)^{k-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd}))+\sum_{i=0}^{k+1}\eta(1-\eta)^{k+1-i}{\bm{w}}_{i}$. Assumption 3 ensures the existence of the gradient and the Hessian of $f_{\theta}$. We have,
$f_{\theta}({\bm{x}}_{k+1}^{sgd})\approx f_{\theta}({\bm{x}}_{k+1}^{\prime})+\bm{\delta}_{k+1}^{\top}\nabla_{\bm{x}}f_{\theta}({\bm{x}}_{k+1}^{\prime})+\frac{1}{2}\bm{\delta}_{k+1}^{\top}\nabla^{2}_{\bm{x}}f_{\theta}({\bm{x}}_{k+1}^{\prime})\bm{\delta}_{k+1}.$
The higher the iteration $k$, the larger penalty weight $\eta(1-\eta)^{K-1-i}$ becomes. Thus with a large $K$, SGD jittering promotes a smoother and wider landscape for the iterative inputs of the $f_{\theta}$. We also refer to this as a flat minimum because the update of ${\bm{x}}_{k}$ solves the lower-level minimization problem. This is consistent with the observations in (Lim et al. 2021), which demonstrate that adding random Gaussian noise to recurrent neural networks layer-wise promotes a small Hessian in the last single-block neural network in terms of its inputs. While it is widely observed that flat minima in the loss landscape (with respect to network parameters) imply better generalization as summarized in Related Works, (Lim et al. 2021) does not clarify why flat minima with respect to hidden-layer inputs improve generalization in regression settings. Our work fills this gap in the context of solving IPs. Now, we state the robustness result,
Theorem 3.
Let ${\bm{e}}$ be the perturbation vector sampled i.i.d. from the zero-mean Gaussian distribution, $\mathcal{P}_{e}$, such that $\mathbb{E}||{\bm{e}}||_{2}^{2}=\sigma_{\bm{e}}^{2}$. For an inverse mapping $H_{\theta}$, the average-case robustness risk for denoising is as follows,
$R_{e}(\theta)=\mathbb{E}_{({\bm{x}},{\bm{y}})\in\mathcal{D},{\bm{e}}\in\mathcal{P}_{e}}~{}\left[||{\bm{x}}-H_{\theta}({\bm{y}}+{\bm{e}})||^{2}_{2}\right].$ (9)
SGD jittering is more robust than MSE training against bounded-variance perturbations around the measurements,
$R_{e}(\theta^{sgd})\leq R_{e}(\theta^{mse}).$
The proof is similar to the one in proving generalization, we include the full proof in the Appendix. The gradient and Hessian of $f_{\theta}$ together help characterize the generalization and robustness of MBAs.
Figure 2: Seismic deconvolution under adversarial attack.
Experiments and Discussion
In our experiments, we use a toy denoising example to validate our theorems and demonstrate the proposed training scheme using seismic deconvolution and magnetic resonance imaging (MRI) reconstruction. For each task, we train the same LU architecture using different training schemes. We then evaluate the models using in-distribution testing data, adversarial attacks with various strengths, and task-dependent OOD data. Detailed training procedures are provided in the Appendix.
Toy problem: 2-dimensional Point Denoising
To validate the Theorems, we compare the robustness and generalization using regular MSE, AT and SGD-jittering in a point denoising problem. Assume ${\bm{x}}=(0,0)^{\top}$, we generate noisy ${\bm{y}}$ by adding random zero-mean Gaussian noise with a variance of 0.01. We use 200 samples for training and 50 for evaluation. Figure 1 shows the noisy points ${\bm{y}}$ and the denoised results $\hat{{\bm{x}}}$ (blue) across four evaluation modes, with the ground-truth in red. In-distribution data are sampled from the same distribution as the training data, centered at $(0,0)^{\top}$. The worst-case attack finds a vector within a 0.01 $\ell_{2}$ distance around ${\bm{y}}$ that maximizes the squared error, while the average-case attack uniformly samples attack vectors within the same ball and adds to ${\bm{y}}$. SGD jittering and AT are more robust than MSE training in both metrics. AT aims to promote worst-case robustness, which implies average-case robustness for attacks within the same $\ell_{2}$ ball around ${\bm{y}}$, though its average-case performance may not be optimal. We observe that SGD jittering has even tighter clusters in its reconstructions under average-case attacks than AT. Finally, we evaluate generalization accuracy by reconstructing the slightly biased ground truth $(0.005,0)^{\top}$ from its corresponding measurements. With the introduction of bias, SGD jittering achieves a more accurate reconstruction compared to the new ground truth. Notice that MSE training underperforms on in-distribution data compared to SGD jittering, likely because the noise injection in SGD jittering helps the model escape from local minima in training.
Figure 3: Seismic deconvolution with an additional layer.
Figure 4: MRI evaluation (a) under adversarial attack obtained from projected gradient descent with norm no more than $\epsilon=1$. Each row represents an MRI reconstruction of a sample from fastMRI (Zbontar et al. 2018a). (b) MRI evaluation using OOD tumor knee data from (Bickle and Jin 2021).
Seismic Deconvolution
Seismic deconvolution is a crucial problem in geophysics. It aims to reverse the convolution effects in the received signal, which occurs when artificial source waves travel through the Earth’s layers. The goal is to extract sparse reflectivity series from recorded seismic data. The training data is generated following the same procedure as in (Iqbal et al. 2019; Guan et al. 2024). To demonstrate the generalization accuracy, we generate OOD testing data by adding an arbitrary horizontal layer with magnitudes of 0.1 and 0.3 to the ground truth, reflected accordingly in the measurements. Table 1 numerically compares the average testing peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) of the reconstructions. Notice that for OOD data, the numerical values are very similar, as the additional layer has a small magnitude relative to the overall measurement. However, the difference can be clearly observed in the figures. Figure 2 shows two deconvolution examples for each trained model under adversarial attacks. MSE training produces artifacts, while both AT and SGD jittering yield results closer to the ground truth. Notably, in the top example, AT fails to recover the small horizontal layer in the center of the boxed area, but SGD jittering successfully captures it. Figure 3 demonstrates the generalization ability with two different magnitude levels. For the OOD example with the larger magnitude, both SGD jittering and MSE training recover a cleaner layer compared to AT. More notably, with the smaller magnitude, AT perceives the line as an unwanted perturbation and barely recovers it, whereas the result from SGD jittering is more distinctly visible.
PSNR / SSIM Test data Adv. Attack OOD small OOD large
MSE training 34.64 / 0.921 28.85 / 0.829 34.57 / 0.918 34.57 / 0.917
AT 33.50 / 0.903 30.27 / 0.849 33.45 / 0.902 33.43 / 0.902
Input Jittering 32.92 / 0.882 30.10 / 0.832 32.91 / 0.882 32.90 / 0.882
SGD jittering 35.10 / 0.928 29.89 / 0.842 34.93 / 0.927 34.93 / 0.927
Table 1: Seismic deconvolution evaluation for in-distribution data (column 2), under adversarial attack with $\epsilon=1$ (column 3), and for OOD data with small and large magnitudes (columns 4 and 5). The best and second-best performances are in bold and underlined respectively.
Accelerated MRI Reconstruction
Accelerated MRI aims to recover human-interpreted body structures from partial k-space measurements. We train models with single-coil knee MRI from the fastMRI dataset (Zbontar et al. 2018b) with 4×4$\times$4 × acceleration, or $1/4$ of the measurements in k-space is used for reconstruction. To assess generalization, we use a different knee dataset from Bickle and Jin (2021), which includes giant tumor cells not present in the training dataset. Table 2 numerically compares the average testing PSNR and SSIM of the reconstructions for in-distribution fastMRI data, under adversarial attack with $\epsilon=1$ and OOD tumor cell knee data. Our findings indicate that MSE training and SGD jittering exhibit comparable performance on in-distribution data, but SGD jittering demonstrates greater robustness and superior performance on OOD data. Adversarial training (AT) yields the most robust model under attack, though it comes with a significant tradeoff in performance for both in-distribution and OOD data. Additionally, as shown in Figure 4, reconstructions using AT appear blurrier than others, a similar observation noted in black-box IP solvers by Krainovic, Soltanolkotabi, and Heckel (2023).
PSNR / SSIM fastMRI Adv. Attack Tumor Cell
MSE training 28.21 / 0.603 25.68 / 0.382 29.92 / 0.779
AT 27.68 / 0.564 27.17 / 0.549 27.74 / 0.597
Input Jittering 28.18 / 0.595 25.05 / 0.420 29.97 / 0.740
SGD jittering 28.22 / 0.607 26.77 / 0.552 30.36 / 0.788
Table 2: MRI evaluation for in-distribution data (column 2), adversarial attack (column 3), and OOD data (column 4). The best and second best performances are in bold and underlined respectively.
Conclusion
Robustness and generalization accuracy are both crucial in solving IPs, yet current analyses often focus on simple black-box solvers that enhance robustness at the cost of accuracy. While empirical studies have examined the robustness of MBAs, there is a lack of theoretical analysis. In this work, we introduce a novel and easily implemented SGD jittering training scheme to address the tradeoff between robustness and accuracy in solving IPs. Our method proves effective for high-quality model-based IP solvers. Through mathematical analysis, we demonstrate that this approach implicitly regularizes the gradient and Hessian of the neural network with respect to the input, resulting in improved generalization and enhanced robustness compared to traditional MSE training. Experimental results show that while MSE training is vulnerable to adversarial attacks and AT recovers smooth estimates, SGD jittering consistently produces robust, high-quality outcomes.
Appendix A Appendix
Appendix B Notations
We clarify the notations used in the paper.
  • $H_{\theta}$: learned inverse mapping parameterized by $\theta$, i.e., black-box neural networks, MBAs. For MBAs in particular, $H_{\theta}=\operatorname*{arg\,min}_{x}~{}\frac{1}{2}\|{\bm{y}}-\bm{A}{\bm{x}}\|^{2}_{2}+r_{\theta}({\bm{x}})$.
  • $r$: regularization function from predefined functions, i.e., $\ell_{1}$ and $\ell_{2}$ norm.
  • $r_{\theta}$: regularization function whose gradient is learned from a neural network $f_{\theta}$.
  • $f_{\theta}$: learned gradient update of $r$, $f_{\theta}({\bm{x}})=\nabla_{\bm{x}}r_{\theta}({\bm{x}})$.
  • $F(x)=\frac{1}{2}\|{\bm{y}}-\bm{A}{\bm{x}}\|^{2}_{2}+r_{\theta}({\bm{x}})$, denote the value of the lower-level objective function.
Appendix C Proof of Lemma 1
Lemma 4.
(Restate Lemma 1. Convergence of MBAs-SGD for IPs) Let $F({\bm{x}})=\frac{1}{2}||{\bm{y}}-\bm{A}{\bm{x}}||_{2}^{2}+r({\bm{x}})$ denote the lower-level objective function. Assume $f_{\theta}({\bm{x}})$ is $L$-Lipschitz continuous,
$||f_{\theta}({\bm{x}}+\Delta)-f_{\theta}({\bm{x}})||_{2}\leq L||\Delta||_{2},\quad\forall{\bm{x}},\Delta$
With assumption 1, and step-sizes $\eta_{k}<2/L$, SGD in minimizing $F({\bm{x}})$ converges to a stationary point.
Proof.
The lemma can be easily derived from the Theorem in (Ghadimi and Lan 2013), which assumes $\nabla_{\bm{x}}F({\bm{x}})$ is $L$-Lipschitz continuous.
$||\nabla_{\bm{x}}F({\bm{x}}+\Delta)-\nabla_{\bm{x}}F({\bm{x}})||_{2}$
$=$ $||\bm{A}^{\top}(\bm{A}({\bm{x}}+\Delta)-{\bm{y}})+f_{\theta}({\bm{x}}+\Delta)-\bm{A}^{\top}(\bm{A}{\bm{x}}-{\bm{y}})-f_{\theta}({\bm{x}})||_{2}$
$=$ $||\bm{A}^{\top}\bm{A}\Delta+f_{\theta}({\bm{x}}+\Delta)-f_{\theta}({\bm{x}})||_{2}$
$\leq$ $||\bm{A}^{\top}\bm{A}||\,||\Delta||_{2}+||f_{\theta}({\bm{x}}+\Delta)-f_{\theta}({\bm{x}})||_{2}$
$\leq$ $\max_{i=1,..,n}\lambda_{i}||\Delta||_{2}+||f_{\theta}({\bm{x}}+\Delta)-f_{\theta}({\bm{x}})||_{2}$
where $\lambda_{i}$ be the eigenvalues of $\bm{A}^{\top}\bm{A}$, and let $\mu=\max_{i=1,..,n}\lambda_{i}$. Since $\nabla F$ is $L$-Lipshitz, we show that for MBAs-SGD in solving denoising problems, it is equivalent to $(L-\mu)$-Lipschitz continuous of $f_{\theta}$. With the same assumptions on step sizes and the SGD jittering noise, we finish the proof. ∎
Appendix D Proof of Theorem 3 - robustness
Theorem 5.
(Restate Theorem 3) Let ${\bm{e}}$ be the perturbation vector and let $\mathcal{P}_{e}$ be iid zero-mean Gaussian distribution such that $\mathbb{E}||{\bm{e}}||_{2}^{2}=\sigma_{\bm{e}}^{2}$. The average-case robustness risk for denoising problem is as follows,
$R_{e}(\theta)=\mathbb{E}_{({\bm{x}},{\bm{y}})\in\mathcal{D},{\bm{e}}}~{}\left[||{\bm{x}}-H_{\theta}({\bm{y}}+{\bm{e}})||^{2}_{2}\right].$ (10)
For the denoising problem, SGD jittering is more robust than MSE training against bounded-variance perturbations around the measurement, or
$R_{e}(\theta^{sgd})\leq R_{e}(\theta^{mse}).$
Proof.
The proof is similar to proving generalization accuracy. To evaluate the average-case robustness for any trained $\theta$, we first write the iterative updates under attack $\{{\bm{x}}_{0}^{at},{\bm{x}}_{1}^{at},...{\bm{x}}_{K}^{at}\}$ in terms of the noiseless trajectory $\{{\bm{x}}_{0}^{\prime},{\bm{x}}_{1}^{\prime},...{\bm{x}}_{K}^{\prime}\}$.
${\bm{x}}_{0}^{at}$ $={\bm{x}}_{0}^{\prime}+{\bm{e}}={\bm{y}}+{\bm{e}}$
${\bm{x}}_{1}^{at}$ $=(1-\eta){\bm{x}}_{0}^{at}+\eta({\bm{y}}+{\bm{e}})-\eta f_{\theta}({\bm{x}}_{0}^{at})$
$=$ ${\bm{x}}_{1}^{\prime}+{\bm{e}}+\eta(f_{\theta}({\bm{x}}_{0}^{\prime})-f_{\theta}({\bm{x}}_{0}^{at}))$
${\bm{x}}_{2}^{at}$ $=(1-\eta){\bm{x}}_{1}^{at}+\eta({\bm{y}}+{\bm{e}})-\eta f_{\theta}({\bm{x}}_{1}^{at})$
$=$ ${\bm{x}}_{2}^{\prime}+{\bm{e}}+\eta(1-\eta)(f_{\theta}({\bm{x}}_{0}^{\prime})-f_{\theta}({\bm{x}}_{0}^{at}))+\eta(f_{\theta}({\bm{x}}_{1}^{\prime})-f_{\theta}({\bm{x}}_{1}^{at}))$
$\quad$...
${\bm{x}}_{k+1}^{at}$ $={\bm{x}}_{k+1}^{\prime}+{\bm{e}}+\sum_{i=0}^{k}\eta(1-\eta)^{k-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))$
Let $K=k+1$, we rewrite the robustness risk. Notice that the generalization loss in (8) includes the additional vector ${\bm{g}}$, whereas the robustness risk in (9) does not.
$R_{e}(\theta)=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}-{\bm{e}}-\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))||^{2}_{2}$
$=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))||^{2}_{2}$
$\quad+\mathbb{E}||{\bm{e}}||_{2}^{2}+2\mathbb{E}\left[<{\bm{e}},\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))>\right]$
$=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))||^{2}_{2}$
$\quad+\mathbb{E}||{\bm{e}}||_{2}^{2}+2\mathbb{E}\left[<{\bm{e}},-\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}-f_{\theta}({\bm{x}}_{i}^{at})>\right]$
$=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))||^{2}_{2}+\sigma_{\bm{e}}^{2}$
$\quad+2\mathbb{E}\left[\langle{\bm{e}},-\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}-f_{\theta}({\bm{x}}_{i}^{at})\rangle\right]$
The first term is the mean-squared testing loss for in-distribution data. The second term is an additional penalty that appears in SGD jittering training. When picking $\sigma_{{\bm{w}}_{k}}^{2}$ for all $k$ and $\sigma_{\bm{e}}^{2}$ such that $\sum_{k=0}^{k}\eta^{2}(1-\eta)^{2(k-i)}\sigma_{w_{i}}^{2}=\sigma_{\bm{e}}^{2}$, this term in robustness risk is a special case in SGD jittering risk. The last term also appears in SGD jittering risk, when the above condition for jittering level and robustness level are satisfied, ${\bm{e}}$ and $\sum_{i=0}^{K}\eta(1-\eta)^{K-i}{\bm{w}}_{i}$ are from the same distribution. Therefore, for models that perform equally well in mean-squared error for in-distribution testing data, minimizing SGD jittering risk minimizes overall robustness. ∎
Appendix E Training Details
We use 10-iteration gradient descent loop unrolling architectures for all tasks. For the toy example, we use a 3-layer MLP with a hidden dimension of 32 as the learned gradient network. For seismic deconvolution and MRI reconstruction, a 5-layer and 8-layer DnCNN with 64 hidden channels are used for the learned gradient network, respectively. All models are trained using Nvidia RTX3080, using Adam optimizer with a learning rate of $41e-$ 4.
The jittering noise variance for each method is selected based on performances of both robustness and accuracy. In input jittering training, the three methods in order of 2D denoising, seismic deconvolution and MRI reconstruction use the noise variance of 0.01, 0.05 and 0.05. In SGD jittering training, we choose the DGD jittering variance for each task with 0.01, 0.1 and 0.01 respectively.
Appendix F Training Speed
We also compare the training speed among different training schemes. All methods are trained on the same device. Training batch sizes for the 2D denoising problem, seismic deconvolution and MRI are 256, 16, and 4 respectively. AT is trained by projected gradient descent. For the 2D denoising problem, $\epsilon=0.01$ with 50 PGD iterations and a stepsize of 0.05. For the seismic deconvolution and MRI reconstruction problem, $\epsilon=1$, with 20 PGD iterations and a stepsize of 0.1.
2D denoising Seis. Deconv. MRI
MSE training 5360.84 it/s 23.69 it/s 5.23 it/s
AT 562.75 it/s 1.53 it/s 0.25 it/s
Input Jittering 5256.77 it/s 22.80 it/s 5.20 it/s
SGD jittering 4337.84 it/s 22.72 it/s 5.27 it/s
Table 3: Training speed measured in items per second are recorded.
AT is significantly slower than other methods. Notice that SGD jittering is very time efficient, even with the additional noise sampling steps.

### v2

@@
-Stochastic Gradient Descent Jittering for Inverse Problems: Alleviating the Accuracy-Robustness Tradeoff
+SGD Jittering: A Training Strategy for Robust and Accurate Model-Based Architectures
 Abstract
@@
 We validate SGD jittering using denoising toy examples, seismic deconvolution, and single-coil MRI reconstruction.
-The proposed method achieves cleaner reconstructions for out-of-distribution data and demonstrates enhanced robustness to adversarial attacks.
-Introduction
+Both SGD jittering and its SPGD extension yield cleaner reconstructions for out-of-distribution data and demonstrates enhanced robustness against adversarial attacks.
+Machine Learning, ICML
+1 Introduction
 Inverse problems (IPs) aim to recover the underlying signal ${\bm{x}}\in\mathbb{R}^{n}$ from corrupted observation ${\bm{y}}\in\mathbb{R}^{m}$, where $m$ is often less than $n$, leading to non-unique solutions.
@@
 where $\bm{A}\in\mathbb{R}^{m\times n}$ is a known forward model for linear IPs, and ${\bm{z}}$ represents unknown noise in the observation.
-Machine learning (ML) algorithms show significantly better reconstructions than the classical optimization methods.
-The majority of the work in the field aims to promote reconstruction quality using different deep-learning models, such as transformers (Fabian, Tinaz, and Soltanolkotabi 2022) and diffusion models (Whang et al.
-2022).
+Machine learning algorithms show significantly better reconstructions than classical optimization methods.
+Most of the work in the field aims to promote reconstruction quality using different deep-learning models, such as transformers (Fabian et al., 2022) and diffusion models (Whang et al., 2022).
 In particular, model-based architectures (MBAs) such as loop unrolling (LU) networks inspired by classical optimization methods, unroll the objective function into a sequence of trainable iterative steps with the forward model $\bm{A}$ involved in every iteration.
-MBAs are considered to be more interpretable than neural networks that learn a direct inverse without using $\bm{A}$ (Monga, Li, and Eldar 2021), and achieve state-of-the-art reconstructions in different fields, such as natural image restoration (Gilton, Ongie, and Willett 2021) and medical image reconstruction (Bian 2024).
+MBAs are considered to be more interpretable than neural networks that learn a direct inverse without using $\bm{A}$ (Monga et al., 2021), and achieve state-of-the-art reconstructions in different fields, such as natural image restoration (Gilton et al., 2021) and medical image reconstruction (Bian, 2024).
 While reconstruction quality is crucial, robustness and generalization accuracy are equally important for IPs.
@@
 Despite their importance, robustness and generalization have not been thoroughly explored in the context of IPs, particularly in MBAs.
-While some studies provide empirical assessments of robustness (Gilton, Ongie, and Willett 2021; Genzel, Macdonald, and März 2022), there is a noticeable lack of theoretical analysis (Gottschling et al.
-2020).
+While some studies provide empirical assessments of robustness (Gilton et al., 2021; Genzel et al., 2022), there is a noticeable lack of theoretical analysis (Gottschling et al., 2025).
 In this work, we model MBA training as a bilevel optimization problem without making specific assumptions about neural network architectures.
-We theoretically analyze its generalization and robustness risks in solving denoising problems and demonstrate that standard mean-squared error (MSE) training may perform poorly in both metrics.
-Technologies like adversarial training (AT) (Fawzi, Moosavi-Dezfooli, and Frossard 2016; Ford et al.
-2019; Kannan, Kurakin, and Goodfellow 2018) and input noise injection (Fawzi, Moosavi-Dezfooli, and Frossard 2016; Ford et al.
-2019; Kannan, Kurakin, and Goodfellow 2018) are commonly used to enhance robustness, while approaches such as weight perturbation (Wu, Xia, and Wang 2020), stochastic training (Hardt, Recht, and Singer 2016) and sharpness-aware minimization (Foret et al.
-2020) are designed to improve generalization.
-However, there is often a tradeoff between robustness and accuracy (Tsipras et al.
-2018; Zhang et al.
-2023; Gottschling et al.
-2020; Krainovic, Soltanolkotabi, and Heckel 2023).
-Although strategies like careful loss function design (Pang et al.
-2022; Kannan, Kurakin, and Goodfellow 2018; Zhang et al.
-2019; Zheng, Zhang, and Mao 2021), data augmentation (Zhang et al.
-2017), and robust architectural choices such as dropout (Yang et al.
-2020) and teacher-student models (Zhao, Wang, and Wei 2024) have been explored to mitigate this tradeoff in classification tasks, the challenge of balancing accuracy and robustness in IP solvers still remains an open question (Gottschling et al.
-2020).
-Recent work by Krainovic, Soltanolkotabi, and Heckel (2023) introduces jittering in ${\bm{y}}$ for black-box IP neural networks, achieving a comparable level of worst-case robustness to AT but with significant accuracy loss.
+We theoretically analyze its generalization and robustness risks in solving denoising problems and demonstrate that standard mean squared error (MSE) training may perform poorly in both metrics.
+Technologies like adversarial training (AT) (Fawzi et al., 2016; Gilmer et al., 2019; Kannan et al., 2018) and input noise injection (Fawzi et al., 2016; Gilmer et al., 2019; Kannan et al., 2018) are commonly used to enhance robustness, while approaches such as weight perturbation (Wu et al., 2020), stochastic training (Hardt et al., 2016) and sharpness-aware minimization (Foret et al., 2021) are designed to improve generalization.
+However, there is often a tradeoff between robustness and accuracy (Tsipras et al., 2019; Zhang et al., 2023; Gottschling et al., 2025; Krainovic et al., 2023).
+Strategies like careful loss function design (Pang et al., 2022; Kannan et al., 2018; Zhang et al., 2019; Zheng et al., 2021), data augmentation (Zhang et al., 2018), and robust architectural choices such as dropout (Yang et al., 2020) and teacher-student models (Zhao et al., 2024) have been explored to mitigate this tradeoff in classification tasks, but the challenge of balancing accuracy and robustness in IP solvers still remains an open question (Gottschling et al., 2025).
+Recent work by Krainovic et al.
+(2023) introduces jittering in ${\bm{y}}$ for black-box IP neural networks, achieving a comparable level of worst-case robustness to AT but with significant accuracy loss.
 To bridge this gap, this work aims to improve robustness through a novel training scheme for MBAs, while maintaining generalization accuracy for IPs.
@@
 Our contribution can be summarized as follows,
-  • (network assumptions) We investigate the performance of MBAs with general non-convex neural networks, where training often results in non-unique parameters.
+  • Network Assumptions: We investigate the performance of MBAs with general non-convex neural networks, where training often results in non-unique parameters.
 This analysis is crucial for demonstrating improvements across various metrics, particularly in robustness and generalization accuracy.
-  • (robustness) We theoretically prove that SGD jittering implicitly promotes average-case robustness than regular MSE training.
+  • Robustness: We theoretically prove that SGD jittering implicitly promotes average-case robustness than regular MSE training.
 In the experiments, we further demonstrate the effectiveness of SGD jittering under worst-case attacks.
-  • (generalization) We are the first to analyze generalization accuracy for IPs with small perturbations in the testing data, rather than relying on PAC-Bayes Theorem, which hypothetically promotes generalization through Hessian regularization.
-  • (convergence) We also demonstrate that under the assumption of Lipschitz continuity, SGD jittering converges, ensuring the training is feasible.
-Related Work
+  • Generalization: To the best of our knowledge, this is the first work to formally analyze generalization accuracy in inverse problems under small perturbations in the test data, by reformulating the generalization risk specific to inverse problems.
+This goes beyond the commonly observed heuristic that a flatter loss landscape implies better generalization.
+2 Related Work
+Robustness and Generalization for IPs
+Several strategies have been proposed to improve robustness in solving inverse problems, including training-time exposure to diverse perturbations such as noise injection and data augmentation (Krainovic et al., 2023; Zhou et al., 2019), and adversarial training (Calivá et al., 2021), enabling models to better handle noise during inference.
+Other approaches focus on architectural innovations—for example, diffusion models have shown inherent robustness to noise in MRI reconstruction (Güngör et al., 2023), while PINN (Peng et al., 2022) embed domain knowledge to enhance stability without compromising interpretability.
+To improve generalization in IPs, methods like data augmentation with synthetic perturbations (Guan et al., 2024) and domain adaptation via techniques such as CycleGAN (Zhu et al., 2017) have been used to expand the training distribution and improve adaptability to unseen data.
+Additionally, incorporating geometric constraints, as in (Jiang et al., 2022) for electrocardiographic image reconstruction, has shown improved generalization by embedding prior knowledge into the learning process.
+While these approaches typically target either robustness or generalization, achieving both simultaneously remains an open and challenging problem.
 Analysis of Robustness
@@
 Works like Salman et al.
-(2019); Cohen, Rosenfeld, and Kolter (2019); Lecuyer et al.
+(2019); Cohen et al.
+(2019); Lecuyer et al.
 (2019); Li et al.
@@
 On the other hand, the analysis of IP is very different from classifications.
-Krainovic, Soltanolkotabi, and Heckel (2023) analyzes the worst-case robustness for IPs using a simple linear reconstruction model and directly solves for robustness risk.
+Krainovic et al.
+(2023) analyzes the worst-case robustness for IPs using a simple linear reconstruction model and directly solves for robustness risk.
 However, optimizing a convex model results in a solution that excels in only one metric (accuracy or robustness).
-In this work, we examine nonlinear MBAs in general forms to show improvements in both metrics.
+In this work, we explore nonlinear MBAs in their general forms, demonstrating improvements across both metrics.
 Analysis of Generalization
-Demonstrating better generalization of classifiers often involves showing larger classification margins (Cao et al.
-2019; Kawaguchi, Kaelbling, and Bengio 2017; Lim et al.
-2021), but the proof for IPs is very different.
+Demonstrating better generalization of classifiers often involves showing larger classification margins (Cao et al., 2019; Kawaguchi et al., 2022; Lim et al., 2021), but the proof for IPs is very different.
 A common but indirect approach shows an algorithm can regularize the Hessian of loss landscape with respect to the network parameters.
-Models that find flat minima in the loss landscape tend to generalize better in practice (Jiang et al.
-2019; Keskar et al.
-2016; Neyshabur et al.
-2017; Foret et al.
-2020).
-Many works demonstrate regularization of the Hessian (Xie, Sato, and Sugiyama 2020; Keskar et al.
-2016; Foret et al.
-2020) then complete the proof using the PAC-Bayes Theorem in hypothesis.
-While the PAC-Bayes theorem provides a bound on the generalization error, it does not directly incorporate curvature measures.
-The exact theoretical mechanism behind it is still not fully understood.
-In this work, instead of connecting to the PAC-Bayes Theorem, we rewrite the generalization risk and demonstrate that SGD jittering results in a smaller generalization risk than regular MSE training.
+Models that converge to flat minima in the loss landscape tend to generalize better in practice (Jiang* et al., 2020; Keskar et al., 2017; Neyshabur et al., 2017; Foret et al., 2021), although the precise theoretical mechanism underlying this phenomenon remains unclear.
+Many studies have shown that regularizing the Hessian of the loss landscape (Xie et al., 2021; Keskar et al., 2017; Foret et al., 2021) leads to flatter minima and have directly linked this to reduced generalization error, but this connection still lacks comprehensive theoretical justification.
+In this work, rather than focusing on the connection to flat landscapes, we proposed a new generalization risk for IPs, and demonstrate that SGD jittering results in a smaller generalization risk than regular MSE training.
 Noise Injections
-Noise injection can be viewed as implicit regularization during training, which is employed to enhance model robustness and generalization.
-Input injection adds random noises to the signal space, which can be viewed as a data augmentation strategy.
-It is widely studied in classifications (Fawzi, Moosavi-Dezfooli, and Frossard 2016; Ford et al.
-2019; Rusak et al.
-2020), and recently shown to promote robustness in IPs (Krainovic, Soltanolkotabi, and Heckel 2023).
-Weight injection instead perturbed network parameters during training, which is shown to help escape from local minima (Zhu et al.
-2018; Nguyen et al.
-2019) and promote smooth solutions (Orvieto et al.
-2023; Liu et al.
-2021; Orvieto et al.
-2022).
-Layer-wise noise injection (Lim et al.
-2021) in a recurrent network has been shown to generalize better for classifiers.
+Noise injection serves as an implicit regularization technique during training to enhance model robustness and generalization.
+When applied to input data, it can be viewed as a data augmentation strategy.
+This approach has been extensively studied in classification tasks (Fawzi et al., 2016; Gilmer et al., 2019; Rusak et al., 2020) and has recently demonstrated effectiveness in improving robustness for IPs (Krainovic et al., 2023).
+Layer-wise noise injection (Lim et al., 2021) in a recurrent network has been shown to generalize better for classifiers.
+Weight injection perturbs network parameters during training, which is shown to help escape from local minima (Zhu et al., 2019; Nguyen et al., 2019) and promote smooth solutions (Orvieto et al., 2023; Liu et al., 2021; Orvieto et al., 2022).
 Weights can also be perturbed before parameter updates to enhance generalization, as in stochastic gradient descent (SGD).
 We will explore SGD in detail in the next subsection.
+In a similar vein, Renaud et al.
+(2024) adds noise to plug-and-play algorithms during evaluation to enhance in-distribution performance, though it does not address its impact on robustness and generalization accuracy.
 Connection to SGD Training
-SGD training is effective in promoting generalization through implicit regularization (Smith et al.
-2021).
-In large overparameterized models with non-unique global solutions, SGD favors networks with flat minima (Hochreiter and Schmidhuber 1997).
-The implicit regularization properties of SGD have been explored in contexts such as linear regression (Zou et al.
-2021), logistic regression (Ji and Telgarsky 2018), and convolutional networks (Gunasekar et al.
-2018).
+SGD training is effective in promoting generalization through implicit regularization (Smith et al., 2021).
+In large overparameterized models with non-unique global solutions, SGD favors networks with flat minima (Hochreiter & Schmidhuber, 1997).
+The implicit regularization properties of SGD have been explored in contexts such as linear regression (Zou et al., 2021), logistic regression (Ji & Telgarsky, 2019), and convolutional networks (Gunasekar et al., 2018).
 Unlike SGD training, which adds noise to network parameters, the proposed SGD jittering method introduces noise at each iterative update of the reconstructions in the lower-level optimization.
 Despite its name, SGD jittering is structurally similar to layer-wise noise injection within MBAs.
-Definition of Robustness and Generalization
+3 Definition of Robustness and Generalization
 We formally define average-case robustness and generalization risks for a learned inverse mapping $H_{\theta}:{\bm{y}}\mapsto{\bm{x}}$ in IPs.
-Average-case robustness is also known as the “natural” robustness for average performance (Rice et al.
-2021; Hendrycks and Dietterich 2019), which is equally important as in worst-case.
+Average-case robustness is also known as the “natural” robustness for average performance (Rice et al., 2021; Hendrycks & Dietterich, 2019), which is equally important as in worst-case.
 During evaluation, a small vector ${\bm{g}}$ sampled from the distribution $\mathcal{P}_{g}$ is added to a data point ${\bm{x}}\in\mathcal{D}$.
 The new ground-truth becomes ${\bm{x}}_{g}={\bm{x}}+{\bm{g}}$, with the corresponding observation ${\bm{y}}_{g}=\bm{A}{\bm{x}}_{g}+{\bm{z}}={\bm{y}}+\bm{A}{\bm{g}}$.
-Definition 1.
+Definition 3.1.
 The average-case robustness risk of $H_{\theta}$ measures the distance between the unperturbed ${\bm{x}}$ to the recovered signal $\hat{{\bm{x}}}_{g}=H_{\theta}({\bm{y}}_{g})$ with the presence of ${\bm{g}}$,
-$R_{e}(\theta)=\mathbb{E}_{{\bm{x}},{\bm{y}}\in\mathcal{D},{\bm{g}}\in\mathcal{P}_{g}}||{\bm{x}}-H_{\theta}({\bm{y}}_{g})||^{2}_{2}.$ (2)
-Definition 2.
+$R_{e}(\theta)=\mathbb{E}_{{\bm{x}},{\bm{y}}\sim\mathcal{D},{\bm{g}}\sim\mathcal{P}_{g}}||{\bm{x}}-H_{\theta}({\bm{y}}_{g})||^{2}_{2}.$ (2)
+Definition 3.2.
 The generalization risk of $H_{\theta}$ measures the distance between the actual ground-truth ${\bm{x}}_{g}$ to the recovered signal $\hat{{\bm{x}}}_{g}=H_{\theta}({\bm{y}}_{g})$ with the presence of ${\bm{g}}$,
-$\mathcal{G}(\theta)=\mathbb{E}_{{\bm{x}},{\bm{y}}\in\mathcal{D},{\bm{g}}\in\mathcal{P}_{g}}||{\bm{x}}_{g}-H_{\theta}({\bm{y}}_{g})||^{2}_{2}.$ (3)
-Robustness measures how close the reconstruction is compared to the data within the dataset, and the generalization accuracy measures how well the reconstruction obeys the physics of the inverse problem even for out-of-distribution (OOD) data.
+$\mathcal{G}(\theta)=\mathbb{E}_{{\bm{x}},{\bm{y}}\sim\mathcal{D},{\bm{g}}\sim\mathcal{P}_{g}}||{\bm{x}}_{g}-H_{\theta}({\bm{y}}_{g})||^{2}_{2}.$ (3)
+Robustness measures how close the reconstruction is compared to the data within the dataset, and the generalization accuracy measures how well the reconstruction obeys the physics of the IP even for out-of-distribution (OOD) data.
 Defining generalization in regression is not straightforward, but for IPs, we can achieve this using the underlying forward models.
-Typically, a tradeoff between robustness and accuracy exists for classical feedforward neural networks (Zhang et al.
-2019; Yang et al.
-2020; Stutz, Hein, and Schiele 2019).
+Typically, a tradeoff between robustness and accuracy exists for classical feedforward neural networks (Zhang et al., 2019; Yang et al., 2020; Stutz et al., 2019).
 We aim to illustrate that the SGD jittering technique enhances robustness and accuracy for MBAs without incorporating explicit regularization.
-Model-based Architecture Overview
-MBAs are inspired by classical optimizations, where the underlying signal ${\bm{x}}$ can be estimated via,
+4 Model-based Architecture Overview
+MBAs draw inspiration from classical optimization techniques, where the underlying signal ${\bm{x}}$ is estimated via,
 $\hat{{\bm{x}}}=\operatorname*{arg\,min}_{\bm{x}}\frac{1}{2}||{\bm{y}}-\bm{A}{\bm{x}}||_{2}^{2}+r({\bm{x}}),$
 where $r$ is an arbitrary regularization function to ensure the uniqueness of the reconstruction.
-When $r$ is differentiable, the optimal solution can be solved via gradient descent (GD), iterating through steps $k=1,2,3...$ with a constant step-size $\eta$,
-${\bm{x}}_{k+1}={\bm{x}}_{k}-\eta\bm{A}^{\top}(\bm{A}{\bm{x}}_{k}-{\bm{y}})+\eta\nabla r({\bm{x}}_{k}).$
-In a loop unrolling (LU) architecture using GD, $\nabla r$ is replaced by a shared-weight neural network, denoted by $f_{\theta}$.
-The process begins with an initial estimate ${\bm{x}}_{0}$, which is iteratively updated with a finite number of iterations.
-The final output ${\bm{x}}_{K}$ represents the predicted reconstruction and is compared to the desired target.
-Different variants of LU are summarized in Monga, Li, and Eldar (2021).
-MBA Training Scheme Overview
-In this section, we summarize some common training schemes for general regression problems, adapted here for MBA training.
-Following this, we introduce the proposed SGD jittering method for training MBAs.
+When $r$ is differentiable, the optimal solution can be solved via gradient descent (GD) algorithm, iterating through steps $k=1,2,3...$ with a constant step-size $\eta$,
+${\bm{x}}_{k+1}={\bm{x}}_{k}-\eta\bm{A}^{\top}(\bm{A}{\bm{x}}_{k}-{\bm{y}})-\eta\nabla r({\bm{x}}_{k}).$ (4)
+In a loop unrolling (LU) architecture using GD, the gradient $\nabla r$ is replaced by a neural network, denoted by $f_{\theta}$.
+The process begins with an initial estimate ${\bm{x}}_{0}$, which is iteratively updated using (4) with a fixed number of iterations, denoted by $K$.
+The final output ${\bm{x}}_{K}$ represents the predicted reconstruction and is compared to the desired ground-truth.
+The weights $\theta$ are then updated through end-to-end backpropagation.
+Different variants of LU are summarized in Monga et al.
+(2021).
+5 MBA Training Scheme Overview
+In this section, we summarize some common training schemes for general regression problems, adapted here for MBA training. Following this, we introduce the proposed SGD jittering method.
 Regular MSE Training
-MSE training is the most commonly used training scheme for training regression problems and MBAs, where the goal using MBAs is to minimize the following risk w.r.t. $\theta$,
-$R(\theta)$ $=\mathbb{E}_{({\bm{x}},{\bm{y}})\in\mathcal{D}}||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\quad\textrm{ where, }$, (4)
+Mean-squared error (MSE) is the most commonly used training loss for regression and IPs. In particular, the goal of training a MBA is to minimize the following risk w.r.t. $\theta$,
+$R(\theta)$ $=\mathbb{E}_{{\bm{x}},{\bm{y}}\sim\mathcal{D}}||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\quad\textrm{ where, }$, (5)
 $\hat{{\bm{x}}}$ $=\operatorname*{arg\,min}_{x}~{}\frac{1}{2}\|{\bm{y}}-\bm{A}{\bm{x}}\|^{2}_{2}+r_{\theta}({\bm{x}})$
-Notice that MBAs do to learn $r$ directly, but instead learn $\nabla r$ using a neural network $f_{\theta}$.
+Notice that MBAs do not directly learn $r$, but instead learn its gradient $\nabla r$ using a neural network $f_{\theta}$.
 We denote $r_{\theta}$ to highlight its dependency of $\theta$.
-The goal of MSE training is to minimize the squared distance between the output ${\bm{x}}_{K}$ and the ground-truth ${\bm{x}}$.
-MSE training might overfit the patterns in training data, making it more vulnerable to variations in ${\bm{y}}$.
+The lower-level optimization is implemented using iterative steps, where each iteration corresponding to a single module in a MBA.
+For LU architecture, with a finite and fixed number of iterations $K$, MSE training minimizes the squared distance between the output ${\bm{x}}_{K}$ and the ground-truth ${\bm{x}}$.
+However, this training scheme can overfit to patterns in training data, potentially making the model more vulnerable to variations in ${\bm{y}}$.
 Adversarial Training
-AT is widely recognized as one of the most effective methods for training robust neural networks (Wang et al.
-2021).
+AT is widely recognized as one of the most effective methods for training robust neural networks (Wang et al., 2019).
 AT finds the worst-case attack for each training instance and adds it to the input data designed to deceive the model.
 By learning from these challenging examples, AT improves the model’s robustness.
-AT aims to minimize the worst-case robustness risk for MBAs as follows,
-$R_{\epsilon}(\theta)=\mathbb{E}_{({\bm{x}},{\bm{y}})\in\mathcal{D}}\left[\begin{aligned} &\max_{{\bm{e}}:||{\bm{e}}||_{2}\leq\epsilon}\quad||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\\ &\textrm{s.t.
+For MBA training, AT aims to minimize the worst-case robustness risk as follows,
+$R_{\epsilon}(\theta)=\mathbb{E}_{{\bm{x}},{\bm{y}}\sim\mathcal{D}}\left[\begin{aligned} &\max_{{\bm{e}}:||{\bm{e}}||_{2}\leq\epsilon}\quad||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\\ &\textrm{s.t.
 }\hat{{\bm{x}}}=\operatorname*{arg\,min}_{x}~{}\frac{1}{2}\|{\bm{y}}+{\bm{e}}-\bm{A}{\bm{x}}\|^{2}_{2}+r_{\theta}({\bm{x}})\end{aligned}\right].$ (6)
-Fast gradient sign method (FGSM) and projected gradient descent (PGD) are methods to find the attack vector ${\bm{e}}$, where (Athalye, Carlini, and Wagner 2018) demonstrate that training a defense model with PGD adversarial attacks is more robust.
-However, AT is slow as it requires iterative solvers for the attack vector (Shafahi et al.
-2019).
-AT might suffer from bad generalizations.
-For example, in noiseless IPs, ${\bm{y}}=\bm{A}{\bm{x}}$, the true inverse $\bm{A}^{-1}$ exists.
-AT learns an inverse mapping $H_{\theta}$ so that $H_{\theta}({\bm{y}}+{\bm{e}})={\bm{x}}$, but the true estimated should be $\bm{A}^{-1}({\bm{y}}+{\bm{e}})\neq{\bm{x}}$.
+Fast gradient sign method and projected gradient descent (ProjGD) are common methods to find the attack vector ${\bm{e}}$, where Athalye et al.
+(2018) demonstrates that ProjGD adversarial attacks are more robust.
+However, AT is slow as it requires iterative solvers for the attack vector (Shafahi et al., 2019).
+AT might suffer from poor generalizations.
+Imagine in noiseless IPs where ${\bm{y}}=\bm{A}{\bm{x}}$, and assume the inverse $\bm{A}^{-1}$ exists.
+The goal of AT is to learn an inverse mapping $H_{\theta}$ such that $H_{\theta}({\bm{y}}+{\bm{e}})$ is close to ${\bm{x}}$, for non-zero vector ${\bm{e}}$.
+However, the true estimated is $\bm{A}^{-1}({\bm{y}}+{\bm{e}})={\bm{x}}+\bm{A}^{-1}{\bm{e}}$.
+AT learns to ignore the latter term, thus introducing errors that can degrade generalization accuracy.
 Input Jittering
-Noise injection to the input is widely used in general ML tasks to promote robustness (Salman et al.
-2019; Cohen, Rosenfeld, and Kolter 2019; Lecuyer et al.
-2019).
-In IPs that aim to map from ${\bm{y}}$ to ${\bm{x}}$, (Krainovic, Soltanolkotabi, and Heckel 2023) proposes to add zero-mean random Gaussian vector ${\bm{w}}$ in ${\bm{y}}$ to a neural network that learns a direct reconstruction, and demonstrate that with careful selection of the jittering variance, this training scheme can attain a comparable level of worst-case robustness to that achieved by AT.
-In MBAs, the input jittering approach is expressed as,
-$J^{I}_{\sigma_{w}}(\theta)$ $=\mathbb{E}_{{\bm{w}},({\bm{x}},{\bm{y}})\in\mathcal{D}}||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\quad\textrm{ where,}$, (6)
+Noise injection to the input is widely used in general ML tasks to promote robustness (Salman et al., 2019; Cohen et al., 2019; Lecuyer et al., 2019).
+In IPs that aim to map from ${\bm{y}}$ to ${\bm{x}}$, Krainovic et al.
+(2023) proposes adding a zero-mean random Gaussian vector ${\bm{w}}$ to ${\bm{y}}$ as input to a neural network that learns a direct reconstruction.
+They demonstrate that, with careful selection of the jittering variance, this training scheme can attain a comparable level of worst-case robustness to that achieved by AT.
+In MBAs, the input jittering approach minimizes,
+$J^{I}_{\sigma_{w}}(\theta)$ $=\mathbb{E}_{{\bm{w}},({\bm{x}},{\bm{y}})\sim\mathcal{D}}||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\quad\textrm{ where,}$, (7)
 $\hat{{\bm{x}}}$ $=\operatorname*{arg\,min}_{x}~{}\frac{1}{2}\|{\bm{y}}+{\bm{w}}-\bm{A}{\bm{x}}\|^{2}_{2}+r_{\theta}({\bm{x}}).$
 This method learns an inverse mapping from ${\bm{y}}+{\bm{w}}$ to ${\bm{x}}$, which might suffer from generalization issues in the same way as AT.
-In fact, Krainovic, Soltanolkotabi, and Heckel (2023) demonstrates a significant accuracy loss using input jittering for black-box neural networks.
-Proposed Training Scheme
-Another line of work injects random noises layer-wise into the neural network, where (Hodgkinson et al.
-2021; Jim, Giles, and Horne 1996; Liu et al.
-2020) model the layer-wise injection as stochastic differential equation and show implicit regularization in robustness, and (Lim et al.
-2021) added noises to the recurrent neural network and demonstrates better generalization for classification tasks.
-In this work, we propose adding random noise vectors ${\bm{w}}_{k}$ drawn from a zero-mean Gaussian distribution to the learned gradient updates $f_{\theta}({\bm{x}}_{k})$, with the noise being re-sampled independently for each iteration $k=1,2,3,...$.
-The noisy gradients are unbiased, $\mathbb{E}[f_{\theta}({\bm{x}})+{\bm{w}}_{k}]=\mathbb{E}[f_{\theta}({\bm{x}})]$ for all ${\bm{x}}$.
+In fact, Krainovic et al.
+(2023) demonstrates a significant accuracy loss using input jittering for black-box neural networks.
+6 Proposed Training Scheme
+Another line of work injects noises layer-wise into the neural network, where Hodgkinson et al.
+(2021); Jim et al.
+(1996); Liu et al.
+(2020) model this layer-wise injection as a stochastic differential equation, demonstrating its role in implicit regularization for robustness.
+Additionally, Lim et al.
+(2021) introduces noise into recurrent neural networks and shows improved generalization for classification tasks.
+In this work, we propose adding random zero-mean Gaussian noises, ${\bm{w}}_{k}$, to the gradient updates $f_{\theta}({\bm{x}}_{k})$, with the noise re-sampled independently for each iteration $k=1,2,3,...$.
+The noisy gradients are unbiased, i.e., $\mathbb{E}[f_{\theta}({\bm{x}})+{\bm{w}}_{k}]=\mathbb{E}[f_{\theta}({\bm{x}})]$ for all ${\bm{x}}$.
 Let $\bar{{\bm{w}}}=\{{\bm{w}}_{1},...,{\bm{w}}_{K}\}$.
-The estimated reconstruction is solved iteratively using this noisy gradient descent which is essentially the stochastic gradient descent (SGD) with step size $\eta$.
-Thus, the learning objective is as follows,
-$J^{SGD}_{\sigma_{w_{k}}}(\theta)$ $=\mathbb{E}_{\bar{{\bm{w}}},({\bm{x}},{\bm{y}})\in\mathcal{D}}||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\quad\textrm{ where,}$, (7)
-$\hat{{\bm{x}}}$ $=\lim_{k\rightarrow\infty}{\bm{x}}_{k}-\eta\big{(}\bm{A}^{\top}(\bm{A}{\bm{x}}_{k}-{\bm{y}})+f_{\theta}({\bm{x}}_{k})+{\bm{w}}_{k}\big{)}$
+The estimated reconstruction is solved iteratively using this noisy gradient descent, which is essentially stochastic gradient descent (SGD).
+Thus, with step size $\eta$, the learning objective to minimize the following risk,
+$J^{SGD}_{\sigma_{w_{k}}}(\theta)$ $=\mathbb{E}_{\bar{{\bm{w}}},({\bm{x}},{\bm{y}})\sim\mathcal{D}}||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\quad\textrm{ where,}$, (8)
+$\hat{{\bm{x}}}$ $=\lim_{k\rightarrow\infty}{\bm{x}}_{k}-\eta\big{(}\bm{A}^{\top}(\bm{A}{\bm{x}}_{k}-{\bm{y}})+f_{\theta}({\bm{x}}_{k})+{\bm{w}}_{k}\big{)}.$
+The jittering noise satisfy the following assumption.
+Assumption 6.1.
+(SGD noises) Assume that at each gradient descent iteration $k=1,2,3,...$, noise ${\bm{w}}_{k}\sim\mathcal{N}(0,\sigma_{w_{k}}^{2}/n\mathbf{I})$ is independently sampled from an i.i.d.
+zero-mean Gaussian distribution, where $\mathbb{E}||{\bm{w}}_{k}||_{2}^{2}=\sigma_{w_{k}}^{2}$.
 In AT and input jittering training, the noise vectors are added to ${\bm{y}}$ throughout the entire inversion, introducing a bias to the reconstruction compared to the actual ground-truth.
-In contrast, SGD jittering adds noise to the learned gradient descent updates.
-The lower-level problem in SGD jittering minimizes $\frac{1}{2}||{\bm{y}}-\bm{A}{\bm{x}}||_{2}^{2}+r_{\theta}({\bm{x}})$, with the convergence result in Lemma 1 ensuring the correct inverse mapping even with noise in gradient descent.
-We modified the convergence result of the SGD in (Ghadimi and Lan 2013) for lower-level SGD in MBAs as follows, where the proof of Lemma 1 can be found in the Appendix.
-Assumption 1.
-(SGD noises) Assume noises ${\bm{w}}_{k}\sim\mathcal{N}(0,\sigma_{w_{k}}^{2}/n\mathbf{I})$ sampled from i.i.d.
-zero-mean Gaussian distribution for each $k=1,2,3,...$ are added to each gradient descent iteration, so that $\mathbb{E}||{\bm{w}}_{k}||_{2}^{2}=\sigma_{w_{k}}^{2}$.
-Lemma 1.
-(Convergence of MBAs-SGD for IPs, modified from SGD convergence result in (Ghadimi and Lan 2013)) Let $F({\bm{x}})=\frac{1}{2}||{\bm{y}}-\bm{A}{\bm{x}}||_{2}^{2}+r({\bm{x}})$ denote the lower-level objective function.
-Assume $f_{\theta}({\bm{x}})$ is $L$-Lipschitz continuous,
-$||f_{\theta}({\bm{x}}+\Delta)-f_{\theta}({\bm{x}})||_{2}\leq L||\Delta||_{2},\quad\forall{\bm{x}},\Delta$
-With assumption 1, and step-sizes $\eta_{k}<2/L$, SGD in minimizing $F({\bm{x}})$ converges to a stationary point.
-SGD noise also implicitly enhances robustness.
-Intuitively, exposing a neural network to noisier inputs during training encourages it to converge to smoother regions of the loss landscape, which are less sensitive to input changes.
-This leads to more stable outputs that are less affected by perturbations during evaluation.
-We will analyze how SGD jittering contributes to this robustness in the following section.
-Analysis of Denoising Problem
+In contrast, the lower-level objective for SGD jittering is the same as in regular MSE training.
+Extensive research has explored the convergence properties of SGD in solving nonconvex optimization problems (Ghadimi & Lan, 2013; Li et al., 2024; Madden et al., 2024; Lei et al., 2019), which ensures a correct inverse mapping even with the presence of noise, thus preserves high accuracy in reconstruction.
+We adapted the SGD convergence result in (Garrigos & Gower, 2023) for SGD jittering as follows, where the proof of Corollary 6.2 can be found in the Appendix.
+Corollary 6.2.
+(Convergence of MBAs-SGD for IPs, modified from Theorem 5.12 in (Garrigos & Gower, 2023)) Let $F({\bm{x}})=\frac{1}{2}||{\bm{y}}-\bm{A}{\bm{x}}||_{2}^{2}+r_{\theta}({\bm{x}})$ denote the lower-level objective function in (8).
+Assume $f_{\theta}=\nabla r_{\theta}$ is $L$-Lipschitz continuous, satisfying
+$\footnotesize||f_{\theta}({\bm{x}}+\Delta)-f_{\theta}({\bm{x}})||_{2}\leq L||\Delta||_{2},\quad\forall{\bm{x}},\Delta,$
+Consider a sequence $\{{\bm{x}}_{k}^{sgd}\}_{k=0}^{K}$ generated by SGD in (8) with a constant step-size $\eta=\sqrt{\frac{2}{LL_{\max}K}}$, for $K\geq 1$, the following holds:
+$\min_{0\leq k<K}\mathbb{E}||\nabla F({\bm{x}}_{k}^{sgd})||^{2}_{2}\leq\sqrt{\frac{2LL_{\max}}{K}}\big{(}2(F({\bm{x}}_{0}^{sgd})-\inf F)+\Delta_{F}^{*})\big{)},$
+where, ${\bm{x}}^{*}=\operatorname*{arg\,min}_{x}F({\bm{x}})$, and
+$\Delta_{F}^{*}=F({\bm{x}}^{*})-\inf\big{(}\frac{1}{2}||{\bm{y}}-\bm{A}{\bm{x}}||_{2}^{2}\big{)}-\inf r_{\theta}({\bm{x}}),$
+denote $\lambda_{i}$ as the eigenvalues of $\bm{A}^{\top}\bm{A}$ and $L_{\max}=\max\{L,\max_{i=1,..,n}\lambda_{i}\}$.
+Figure 1: 2D point denoising example using MSE training, AT, and SGD jittering.
+Each method is evaluated using regular in-distribution data, under average- and worst-case attacks to ${\bm{y}}$ and with additional bias (from left to right columns respectively).
+Blue dots represents each test data instance, and the red dots indicate the ground-truth.
+The dashed circle highlights the cluster under attack, and the arrows represent the distance from the center of the data cluster to the biased ground truth.
+On the other hand, SGD jittering noise also implicitly enhances robustness.
+While we will analyze this relationship theoretically in the following section, intuitively, exposing a neural network to noisier inputs during training encourages it to converge to smoother regions of the loss landscape, which are less sensitive to input variations.
+This results in more stable outputs that are less affected by perturbations during evaluation.
+7 Analysis for Denoising Problem
 In this section, we analyze the generalization and robustness of SGD jittering in addressing the denoising problem. We begin by outlining the relevant assumptions.
-Assumption 2.
+Assumption 7.1.
 (Denoising) The forward model in denoising is identity mapping, so that ${\bm{y}}={\bm{x}}+{\bm{z}}$.
@@
 from ${\bm{z}}\sim\mathcal{N}(0,\sigma_{z}^{2}/n\bm{I})$.
-Assumption 3.
-(Neural Network) Assume $f_{\theta}$ is twice differentiable with respect to the input, so the Hessian exists.
-Figure 1: 2D point denoising example using MSE training, AT, and SGD jittering.
-Each method is evaluated using regular in-distribution data, under average- and worst-case attacks to ${\bm{y}}$ and with additional bias (from left to right columns respectively).
-Main Theoretical Results
+Assumption 7.2.
+(Neural Network) Assume the neural network architecture is twice differentiable with respect to the input, so the Hessian exists.
+Definition 7.3.
+Let $\theta^{mse}$, $\theta^{at}$ and $\theta^{sgd}$ denote the optimal neural network parameters of $f$ obtained from regular MSE training in (5), AT in (6) and SGD jittering training in (8), respectively.
+7.1 Main Theoretical Results
 Given the assumptions, we state the first theorem regarding generalization accuracy. Let $H_{\theta}$ denote the inverse process learned using MBAs for the rest of the analysis.
-Theorem 2.
+Theorem 7.4.
 Assuming a small zero-meam Gaussian random vector ${\bm{g}}$ with $\mathbb{E}||{\bm{g}}||^{2}_{2}=\sigma_{g}^{2}$ is added to data ${\bm{x}}\in\mathcal{D}$ during evaluation.
 The generalization risk for denoising problem is defined as,
-$\mathcal{G}(\theta)=\mathbb{E}_{{\bm{x}},{\bm{y}}\in\mathcal{D},{\bm{g}}}||{\bm{x}}+{\bm{g}}-H_{\theta}({\bm{y}}+{\bm{g}})||^{2}_{2}.$ (8)
-Under assumptions 1, 2 and 3, SGD jittering generalizes better than regular MSE training,
+$\footnotesize\mathcal{G}(\theta)=\mathbb{E}_{{\bm{x}},{\bm{y}}\sim\mathcal{D},{\bm{g}}}||{\bm{x}}+{\bm{g}}-H_{\theta}({\bm{y}}+{\bm{g}})||^{2}_{2}.$ (9)
+Under assumptions 6.1, 7.1 and 7.2, SGD jittering generalizes better than regular MSE training,
+$\mathcal{G}(\theta^{sgd})\leq\mathcal{G}(\theta^{mse}).$
+The detailed proof is provided in Appendix C.
+SGD jittering can be interpreted as a noisy version of the regular MSE training.
+The idea is to decompose the generalization accuracy in the presence of ${\bm{g}}$ into two terms: the MSE term corresponding to the reconstruction from the clean trajectory $H_{\theta}({\bm{y}})$, and a penalty term depends on the noisy trajectory $H_{\theta}({\bm{y}}+{\bm{g}})$.
+Then we show that the training loss of SGD jittering includes an implicit regularization term that penalizes deviations in the intermediate reconstructions caused by the perturbation.
+Minimizing the SGD jittering training loss also minimizes the penalty term in $\mathcal{G}(\theta)$, leading to better generalization accuracy compared to regular MSE training.
+Furthermore, we rewrite the regularization term derived in the proof as follows in SGD training via Taylor expansion.
+$\text{regularization}=\mathbb{E}~{}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd}))||_{2}^{2}.$ (10)
+Let $\{{\bm{x}}_{k}^{sgd}\}_{k=0}^{K}$ and $\{{\bm{x}}_{k}^{\prime}\}_{k=0}^{K}$ denote the reconstruction trajectory from SGD and GD, respectively.
+As derived through iterative expansions in the proof, at iteration $k$, we have ${\bm{x}}_{k}^{sgd}={\bm{x}}_{k}^{\prime}+\bm{\delta}_{k}$, where $\bm{\delta}_{k}=\sum_{i=0}^{k-1}\eta(1-\eta)^{k-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd}))+\sum_{i=0}^{k}\eta(1-\eta)^{k-i}{\bm{w}}_{i}$.
+Assumption 7.2 ensures the existence of the first and second derivatives of $f_{\theta}$.
+Let $f_{\theta}^{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be the function defined by $f^{i}_{\theta}({\bm{x}})=[f_{\theta}({\bm{x}})]_{i}$, representing the $i^{th}$ component of $f_{\theta}({\bm{x}})$, for $i\in\{1,...,n\}$.
+Then, we have,
+$f^{i}_{\theta}({\bm{x}}_{k}^{sgd})\approx f^{i}_{\theta}({\bm{x}}_{k}^{\prime})+\bm{\delta}_{k}^{\top}\nabla f_{\theta}^{i}({\bm{x}}_{k}^{\prime})+\frac{1}{2}\bm{\delta}_{k}^{\top}\bm{H}_{f}^{i}({\bm{x}}_{k}^{\prime})\bm{\delta}_{k},$
+where, $\nabla f_{\theta}^{i}$ and $\bm{H}_{f}^{i}$ denote the gradient and the Hessian of $f_{\theta}^{i}$.
+As $k$ increases, the regularization weight $\eta(1-\eta)^{K-1-i}$ in (10) becomes more significant.
+Thus minimizing the SGD jittering risk implicitly promotes smoother and wider landscapes of the lower-level objective $F$ with respect to the iterative inputs ${\bm{x}}_{k}^{\prime}$ for larger values of $K$.
+It is important to note that this differs from the concept of flat minima in a general feedforward neural network, which flatness refers to regions of low training loss with respect to the network parameters, often associated with improved generalization.
+Our findings align with the observations in (Lim et al., 2021), which demonstrate that layer-wise noise injections to recurrent neural networks promotes a smaller Hessian with respect to the hidden inputs to the final layer.
+While it is widely observed that flat minima in the loss landscape (with respect to network parameters) imply better generalization as summarized in Section 2, Lim et al.
+(2021) does not clarify why flat minima with respect to hidden-layer inputs improve generalization in regression settings.
+Our work fills this gap in the context of IPs.
+Furthermore, we present the robustness analysis results.
+Theorem 7.5.
+Let ${\bm{e}}$ be the perturbation vector sampled i.i.d.
+from the zero-mean Gaussian distribution, $\mathcal{P}_{e}$, such that $\mathbb{E}||{\bm{e}}||_{2}^{2}=\sigma_{\bm{e}}^{2}$.
+For an inverse mapping $H_{\theta}$, the average-case robustness risk for denoising is as follows,
+$\footnotesize R_{e}(\theta)=\mathbb{E}_{({\bm{x}},{\bm{y}})\sim\mathcal{D},{\bm{e}}\sim\mathcal{P}_{e}}~{}\left[||{\bm{x}}-H_{\theta}({\bm{y}}+{\bm{e}})||^{2}_{2}\right].$ (11)
+SGD jittering is more robust than MSE training against bounded-variance perturbations around the measurements,
+$R_{e}(\theta^{sgd})\leq R_{e}(\theta^{mse}).$
+The proof is similar to the one in proving generalization, we include the full proof in the Appendix D.
+The implicit regularization induced by SGD jittering help characterize both generalization and robustness of MBAs.
+Figure 2: Visualization of seismic deconvolution using (stochastic) gradient descent variants within loop unrolling architectures.
+Figure (a) illustrates the reconstruction performance under adversarial attack with $\epsilon=1$.
+Figure (b) evaluates the model’s performance on OOD seismic data, where the ground-truth reflectivity includes an additional small horizontal layer (with magnitude of 0.3 and 0.1), highlighted in the red boxes.
+8 Experiments and Discussions
+In our experiments, we use a toy denoising example to validate our theorems and demonstrate the proposed training scheme using seismic deconvolution and magnetic resonance imaging reconstruction.
+For each task, we train the same LU architecture using different training schemes.
+We then evaluate the models using in-distribution (ID) testing data, adversarial attacks with various strengths, and task-dependent OOD data.
+Detailed training procedures are provided in the Appendix.
+Toy Problem: 2-dimensional Point Denoising
+We begin with a 2D point denoising problem to validate our theorems.
+The task involves recovering the point ${\bm{x}}=(0,0)^{\top}$ from noisy observations generated by adding zero-mean Gaussian noise with a variance of 0.01.
+We use 200 samples for training and 50 for evaluation.
+Figure 1 visualizes the denoising results under ID data, average-case attack, worst-case attack, and a biased ground-truth.
+ID data are sampled from the same distribution as the training data, centered at $(0,0)^{\top}$.
+The worst-case attack finds a vector within a 0.01 $\ell_{2}$-distance around ${\bm{y}}$ that maximizes the squared error, while the average-case attack uniformly samples attack vectors within the same ball and adds to ${\bm{y}}$.
+Finally, we evaluate generalization accuracy by reconstructing the slightly biased ground truth $(0.005,0)^{\top}$ from its corresponding measurements.
+The results show that both SGD jittering and AT produce tighter clusters of reconstructed points compared to MSE training under adversarial attacks.
+Moreover, when reconstructing a slightly biased ground truth, SGD jittering achieves the most accurate reconstruction, showcasing its better generalization capability to small variations in data.
+Notice that MSE training underperforms SGD jittering on ID data, likely because the noise injection in SGD jittering helps the model escape from local minima in training.
+PSNR / SSIM ID Test Data Adv. Attack OOD Data
+MSE training 34.64 / 0.921 28.85 / 0.829 34.57 / 0.918
+AT 33.50 / 0.903 30.27 / 0.849 33.45 / 0.902
+Input Jittering 32.92 / 0.882 30.10 / 0.832 32.91 / 0.882
+SGD jittering 35.10 / 0.928 29.89 / 0.842 34.93 / 0.927
+Table 1: Seismic deconvolution evaluation for in-distribution data (column 2), under adversarial attack with $\epsilon=1$ (column 3), and for OOD data with additional horizontal layer (columns 4).
+The best and second-best performances are in bold and underlined respectively.
+Seismic Deconvolution
+Seismic deconvolution is a crucial problem in geophysics.
+It aims to reverse the convolution effects in the received signal, which occurs when artificial source waves travel through the Earth’s layers.
+The goal is to extract sparse reflectivity series from recorded seismic data.
+The training data is generated following the same procedure as in (Iqbal et al., 2019; Guan et al., 2024).
+To test the generalization accuracy, we generate OOD data by introducing an arbitrary horizontal layer with magnitudes of 0.1 to the ground truth, reflected accordingly in the measurements.
+The numerical results in Table 1 show that SGD jittering achieves the highest peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) on ID and OOD data, while maintaining competitive performance under adversarial attacks.
+In contrast, AT demonstrates the best robustness under adversarial attacks but suffers significant performance degradation for both ID and OOD data.
+Visual reconstructions in Figure 2 further highlight the differences between these methods.
+While MSE training often introduces artifacts, especially under adversarial attacks, AT fails to recover the small horizontal layer in OOD data with a magnitude of 0.1, likely perceiving it as noise.
+In comparison, SGD jittering accurately recovers the additional layer, demonstrating its ability to generalize to subtle features not present in the training data.
+PSNR / SSIM fastMRI Adv. Attack Tumor Cell
+MSE training 28.21 / 0.603 25.68 / 0.382 29.92 / 0.779
+AT 27.68 / 0.564 27.17 / 0.549 27.74 / 0.597
+Input Jittering 28.18 / 0.595 25.05 / 0.420 29.97 / 0.740
+SGD jittering 28.22 / 0.607 26.77 / 0.552 30.36 / 0.788
+Table 2: MRI evaluation for in-distribution data (column 2), adversarial attack (column 3), and OOD data (column 4). The best and second best performances are in bold and underlined respectively.
+Figure 3: Gradient descent LU for MRI evaluation (a) under adversarial attack obtained from projected gradient descent with norm no more than $\epsilon=1$.
+Each row represents an MRI reconstruction of a sample from fastMRI (Knoll et al., 2020).
+(b) MRI evaluation using OOD tumor knee data from (Bickle & Jin, 2021).
+Figure 4: Comparing robustness and generalization of different training schemes for GD LU for seismic deconvolution problem (top row), and MRI reconstruction (bottom row).
+Accelerated MRI Reconstruction
+Accelerated MRI aims to recover human-interpreted body structures from partial k-space measurements.
+We train models with single-coil knee MRI from the fastMRI dataset (Knoll et al., 2020) with 4×4$\times$4 × acceleration, or $1/4$ of the measurements in k-space is used for reconstruction.
+To assess generalization, we use a different knee dataset from Bickle & Jin (2021), which includes giant tumor cells absent from the training data.
+The results in Table 2 show that, on ID data, MSE training and SGD jittering perform similarly, while AT underperforms due to its inherent tradeoff in resolution.
+On OOD tumor data, SGD jittering outperforms all other methods, demonstrating better generalization to unseen features.
+Visual comparisons in Figure 3 further support these findings.
+Reconstructions produced by AT are smoother but often miss fine details, consistent with prior observations in black-box IP solvers (Krainovic et al., 2023).
+In contrast, SGD jittering produces high-resolution reconstructions, preserving critical anatomical structures, particularly in OOD cases.
+8.1 Robustness vs. Generalization
+Figure 4 presents a comparison of various training schemes in terms of in-distribution accuracy, robustness and generalization measured by average test PSNR on seismic deconvolution (top row) and MRI reconstruction (bottom row) tasks.
+The left panels show that the proposed SGD jittering approach achieves better generalization to OOD data.
+Meanwhile, the right panels depict the trade-off between robustness and accuracy for baseline methods, where our approach provides a more effective balance between the two metrics.
+8.2 How to Choose SGD Noise Level?
+Figure 6 examines the impact of SGD jittering noise levels $\sigma_{wk}^{2}$ in the context of the denoising problem.
+The jittering noise level determines the robustness and in- and out-of-distribution accuracies.
+This parameter search allows the identification of an optimal $\sigma_{wk}^{2}$ that balances robustness against adversarial attacks while maintaining high accuracy.
+Figure 5: Proximal gradient descent LU for MRI evaluation (a) under adversarial attack obtained from projected gradient descent with norm no more than $\epsilon=1$.
+Each row represents an MRI reconstruction of a sample from fastMRI (Knoll et al., 2020).
+(b) MRI evaluation using OOD tumor knee data from (Bickle & Jin, 2021).
+Figure 6: The effect of SGD jittering noise level $\sigma_{wk}^{2}$ on model performance for denoising problem.
+(Top): Worst-case robustness $R_{\epsilon}$ recorded for each noise level.
+Each line represents a different value of $\epsilon$, with the red dot indicating the $\sigma_{wk}^{2}$ that achieves the best robustness for each corresponding $\epsilon$.
+(Bottom): In-distribution (ID) and out-of-distribution (OOD) accuracies in MSE are recorded for each $\sigma_{wk}^{2}$, with best noise levels highlighted.
+8.3 Stochastic Proximal Gradient Descent Jittering
+We further extend the proposed stochastic training idea to proximal gradient descent (PGD) variant of model-based architecture training, and denote the algorithm as SPGD jittering, where the learning objective is,
+$J^{SPGD}_{\sigma_{w_{k}}}(\theta)$ $=\mathbb{E}_{\bar{{\bm{w}}},({\bm{x}},{\bm{y}})\sim\mathcal{D}}||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\quad\textrm{ where,}$, (12)
+$\hat{{\bm{x}}}$ $=\lim_{k\rightarrow\infty}prox_{\theta}\big{(}{\bm{x}}_{k}-\eta\big{(}\bm{A}^{\top}(\bm{A}{\bm{x}}_{k}-{\bm{y}})+{\bm{w}}_{k}\big{)}\big{)}.$
+Here, $prox$ represents a proximal operator, which is replaced by a neural network parameterized by $\theta$, and ${\bm{w}}_{k}$ denotes the stochastic noises injected independently to the input of the proximal network at each iteration.
+Similar to SGD jittering, SPGD jittering ensures that the noise ${\bm{w}}_{k}$ is re-sampled at every iteration, maintaining unbiased gradient updates while promoting variations in training data.
+Figure 5 compares the same proximal gradient descent framework with different training schemes for accelerated MRI reconstruction.
+MSE training denotes the classical proximal GD variant of LU trained with MSE loss, AT, and SPGD jittering for accelerated MRI reconstruction.
+MSE training demonstrates reasonable generalization to OOD tumor data but remains highly sensitive to adversarial attacks.
+In contrast, AT exhibits robustness against attacks but produces reconstructions with lower resolution, which can obscure fine details in critical medical imaging tasks.
+SPGD jittering effectively enhances both robustness and accuracy, achieving cleaner reconstructions under adversarial attacks and superior generalization to OOD tumor data.
+Numerical evaluations on SPGD jittering for accelerated MRI and further experiments on seismic deconvolution are presented in Appendix E.
+We also compare the training speed of AT and the proposed stochastic noise injection techniques in Appendix H, demonstrating that the proposed method is a time-efficient strategy to promote robustness.
+9 Conclusion
+Robustness and generalization accuracy are both crucial in solving IPs, yet current analyses often focus on simple black-box solvers that enhance robustness at the cost of accuracy.
+While empirical studies have examined the robustness of MBAs, there is a lack of theoretical analysis.
+In this work, we introduce a novel and easily implemented SGD jittering training scheme to address the tradeoff between robustness and accuracy in solving IPs.
+Our method proves effective for high-quality model-based IP solvers.
+Through mathematical analysis, we demonstrate that this approach implicitly regularizes the gradient and Hessian of the neural network with respect to the input, resulting in improved generalization and enhanced robustness compared to traditional MSE training.
+Experimental results show that while MSE training is vulnerable to adversarial attacks and AT recovers smooth estimates, SGD and SPGD jittering consistently produces robust, high-quality outcomes.
+Software and Data
+The code for the proposed SGD and SPGD jittering methods are provided here: https://github.com/InvProbs/SGD-jittering.
+Acknowledgements
+This work was supported by the Center for Energy and Geo Processing (CeGP) at Georgia Institute of Technology.
+Impact Statement
+This work highlights the critical need for both robustness and generalization accuracy in solving inverse problems.
+Model-based architectures are high-performance and data-efficient machine learning architecture for IPs.
+The proposed SGD/SPGD jittering strategy mitigates the robustness-accuracy tradeoff for this specific architecture, demonstrating both efficiency and effectiveness across diverse applications.
+However, for inverse problems with unknown forward models where model-based architectures are unsuitable, future research may require alternative techniques to overcome the robustness-accuracy tradeoff.
+Appendix A Notations
+We clarify the notations used in the paper.
+  • $H_{\theta}$: learned inverse mapping parameterized by $\theta$, i.e., black-box neural networks, MBAs.
+For MBAs in particular, $H_{\theta}=\operatorname*{arg\,min}_{x}~{}\frac{1}{2}\|{\bm{y}}-\bm{A}{\bm{x}}\|^{2}_{2}+r_{\theta}({\bm{x}})$.
+  • $r$: regularization function from predefined functions, i.e., $\ell_{1}$ and $\ell_{2}$ norm.
+  • $r_{\theta}$: regularization function whose gradient is learned from a neural network $f_{\theta}$.
+  • $f_{\theta}$: learned gradient update of $r$, $f_{\theta}({\bm{x}})=\nabla_{\bm{x}}r_{\theta}({\bm{x}})$.
+  • $F(x)=\frac{1}{2}\|{\bm{y}}-\bm{A}{\bm{x}}\|^{2}_{2}+r_{\theta}({\bm{x}})$, denote the value of the lower-level objective function.
+Appendix B Proof of SGD Jittering Convergence (Corollary 6.2)
+Corollary B.1.
+(Restate Corollary 6.2.
+Convergence of MBAs-SGD for IPs) Let $F({\bm{x}})=\frac{1}{2}||{\bm{y}}-\bm{A}{\bm{x}}||_{2}^{2}+r_{\theta}({\bm{x}})$ denote the lower-level objective function.
+Assume $f_{\theta}=\nabla r_{\theta}$ is $L$-Lipschitz continuous, satisfying
+$||f_{\theta}({\bm{x}}+\Delta)-f_{\theta}({\bm{x}})||_{2}\leq L||\Delta||_{2},\quad\forall{\bm{x}},\Delta,$
+let $\mu=\max_{i=1,..,n}\lambda_{i}$, where $\lambda_{i}$ are the eigenvalues of $\bm{A}^{\top}\bm{A}$, and define $L_{max}=\max\{\mu,L\}$.
+Consider a sequence $\{{\bm{x}}_{k}^{sgd}\}_{k=0}^{K}$ generated by SGD in (8) with a constant step-size $\eta=\sqrt{\frac{2}{LL_{max}K}}$.
+For $K\geq 1$, the following holds:
+$\min_{0\leq k<K}\mathbb{E}||\nabla F({\bm{x}}_{k}^{sgd})||^{2}_{2}\leq\sqrt{\frac{2LL_{max}}{K}}\big{(}2(F({\bm{x}}_{0}^{sgd})-\inf F)+\Delta_{F}^{*})\big{)},$
+where, ${\bm{x}}^{*}=\operatorname*{arg\,min}_{x}F$, and
+$\Delta_{F}^{*}=F({\bm{x}}^{*})-\inf\big{(}\frac{1}{2}||{\bm{y}}-\bm{A}{\bm{x}}||_{2}^{2}\big{)}-\inf r_{\theta}({\bm{x}}).$
+Proof.
+The corollary follows directly from Theorem 5.12 in (Garrigos & Gower, 2023), which assumes the objective function is Lipschitz.
+In the case of the SGD jittering training scheme, the lower-level objective function is given by $F({\bm{x}})=\frac{1}{2}||{\bm{y}}-\bm{A}{\bm{x}}||_{2}^{2}+r_{\theta}({\bm{x}})$.
+We show that $F$ is $(L+\mu)$-Lipschitz, allowing the result from Theorem 5.12 to be directly applied.
+Assume $f_{\theta}$ is $L$-Lipschitz, expanding the gradient of $F$, we have,
+$||\nabla_{\bm{x}}F({\bm{x}}+\Delta)-\nabla_{\bm{x}}F({\bm{x}})||_{2}$
+[…]
+[…]
+[…]
+[…]
+where $\lambda_{i}$ be the eigenvalues of $\bm{A}^{\top}\bm{A}$, and let $\mu=\max_{i=1,..,n}\lambda_{i}$. Thus $F$ is $(L+\mu)$-Lipshitz and we conclude the proof.
+Appendix C Proof of Generalization (Theorem 7.4)
+Theorem C.1.
+(Restate Theorem 7.4) Assuming a small zero-meam Gaussian random vector ${\bm{g}}$ with $\mathbb{E}||{\bm{g}}||^{2}_{2}=\sigma_{g}^{2}$ is added to data ${\bm{x}}\in\mathcal{D}$ during evaluation.
+The generalization risk for denoising problem is defined as,
+$\mathcal{G}(\theta)=\mathbb{E}_{{\bm{x}},{\bm{y}}\sim\mathcal{D},{\bm{g}}}~{}||{\bm{x}}+{\bm{g}}-H_{\theta}({\bm{y}}+{\bm{g}})||^{2}_{2}.$ (13)
+Under assumptions 6.1, 7.1 and 7.2, SGD jittering generalizes better than regular MSE training,
 $\mathcal{G}(\theta^{sgd})\leq\mathcal{G}(\theta^{mse}).$
@@
 ${\bm{x}}_{2}^{sgd}$ $=(1-\eta){\bm{x}}_{1}^{sgd}+\eta{\bm{y}}-\eta f_{\theta}({\bm{x}}_{1}^{sgd})-\eta{\bm{w}}_{2}$
-$={\bm{x}}_{2}^{\prime}+\eta\big{(}f_{\theta}({\bm{x}}_{1}^{\prime})-f_{\theta}({\bm{x}}_{1}^{sgd})\big{)}+\eta(1-\eta)\big{(}f_{\theta}({\bm{x}}_{0}^{\prime})-f_{\theta}({\bm{x}}_{0}^{sgd})\big{)}$
-$\qquad-\big{(}\eta(1-\eta)^{2}{\bm{w}}_{0}+\eta(1-\eta){\bm{w}}_{1}+\eta{\bm{w}}_{2}\big{)}$
+$={\bm{x}}_{2}^{\prime}+\eta\big{(}f_{\theta}({\bm{x}}_{1}^{\prime})-f_{\theta}({\bm{x}}_{1}^{sgd})\big{)}+\eta(1-\eta)\big{(}f_{\theta}({\bm{x}}_{0}^{\prime})-f_{\theta}({\bm{x}}_{0}^{sgd})\big{)}-\big{(}\eta(1-\eta)^{2}{\bm{w}}_{0}+\eta(1-\eta){\bm{w}}_{1}+\eta{\bm{w}}_{2}\big{)}$
 ${\bm{x}}_{k+1}^{sgd}$ $=(1-\eta){\bm{x}}_{k}^{sgd}+\eta{\bm{y}}-\eta f_{\theta}({\bm{x}}_{k}^{sgd})-\eta{\bm{w}}_{k+1}$
-$={\bm{x}}_{k+1}^{\prime}+\sum_{i=0}^{k}\eta(1-\eta)^{k-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd}))$
-$\qquad-\sum_{i=0}^{k+1}\eta(1-\eta)^{k+1-i}{\bm{w}}_{i}.$
+$={\bm{x}}_{k+1}^{\prime}+\sum_{i=0}^{k}\eta(1-\eta)^{k-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd}))-\sum_{i=0}^{k+1}\eta(1-\eta)^{k+1-i}{\bm{w}}_{i}.$
 Let $K=k+1$, SGD jittering training risk is,
-$\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{sgd}||_{2}^{2}=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}~{}||\sum_{i=0}^{K}\eta(1-\eta)^{K-i}{\bm{w}}_{i}||_{2}^{2}$
-$+\mathbb{E}~{}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd}))||_{2}^{2}$
-$-2\mathbb{E}\left[\langle\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}f_{\theta}({\bm{x}}_{i}^{sgd}),\sum_{i=0}^{K}\eta(1-\eta)^{K-i}{\bm{w}}_{i}\rangle\right].$
+$\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{sgd}||_{2}^{2}$ $=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}~{}||\sum_{i=0}^{K}\eta(1-\eta)^{K-i}{\bm{w}}_{i}||_{2}^{2}+\mathbb{E}~{}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd}))||_{2}^{2}$
+$\quad+2\mathbb{E}~{}\left[\langle\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd})),\sum_{i=0}^{K}\eta(1-\eta)^{K-i}{\bm{w}}_{i}\rangle\right].$
 The first component corresponds to the MSE loss when the trajectory is not perturbed by jittering noise.
@@
 Therefore, $\mathcal{G}(\theta^{sgd})\leq\mathcal{G}(\theta^{mse})$ due to the implicit regularization term in SGD jittering training.
-In addition, we can expand the regularization term in SGD training via Taylor expansion.
-Let $\bm{\delta}_{k+1}=\sum_{i=0}^{k}\eta(1-\eta)^{k-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd}))+\sum_{i=0}^{k+1}\eta(1-\eta)^{k+1-i}{\bm{w}}_{i}$.
-Assumption 3 ensures the existence of the gradient and the Hessian of $f_{\theta}$.
-We have,
-$f_{\theta}({\bm{x}}_{k+1}^{sgd})\approx f_{\theta}({\bm{x}}_{k+1}^{\prime})+\bm{\delta}_{k+1}^{\top}\nabla_{\bm{x}}f_{\theta}({\bm{x}}_{k+1}^{\prime})+\frac{1}{2}\bm{\delta}_{k+1}^{\top}\nabla^{2}_{\bm{x}}f_{\theta}({\bm{x}}_{k+1}^{\prime})\bm{\delta}_{k+1}.$
-The higher the iteration $k$, the larger penalty weight $\eta(1-\eta)^{K-1-i}$ becomes.
-Thus with a large $K$, SGD jittering promotes a smoother and wider landscape for the iterative inputs of the $f_{\theta}$.
-We also refer to this as a flat minimum because the update of ${\bm{x}}_{k}$ solves the lower-level minimization problem.
-This is consistent with the observations in (Lim et al.
-2021), which demonstrate that adding random Gaussian noise to recurrent neural networks layer-wise promotes a small Hessian in the last single-block neural network in terms of its inputs.
-While it is widely observed that flat minima in the loss landscape (with respect to network parameters) imply better generalization as summarized in Related Works, (Lim et al.
-2021) does not clarify why flat minima with respect to hidden-layer inputs improve generalization in regression settings.
-Our work fills this gap in the context of solving IPs.
-Now, we state the robustness result,
-Theorem 3.
-Let ${\bm{e}}$ be the perturbation vector sampled i.i.d.
-from the zero-mean Gaussian distribution, $\mathcal{P}_{e}$, such that $\mathbb{E}||{\bm{e}}||_{2}^{2}=\sigma_{\bm{e}}^{2}$.
-For an inverse mapping $H_{\theta}$, the average-case robustness risk for denoising is as follows,
-$R_{e}(\theta)=\mathbb{E}_{({\bm{x}},{\bm{y}})\in\mathcal{D},{\bm{e}}\in\mathcal{P}_{e}}~{}\left[||{\bm{x}}-H_{\theta}({\bm{y}}+{\bm{e}})||^{2}_{2}\right].$ (9)
-SGD jittering is more robust than MSE training against bounded-variance perturbations around the measurements,
-$R_{e}(\theta^{sgd})\leq R_{e}(\theta^{mse}).$
-The proof is similar to the one in proving generalization, we include the full proof in the Appendix.
-The gradient and Hessian of $f_{\theta}$ together help characterize the generalization and robustness of MBAs.
-Figure 2: Seismic deconvolution under adversarial attack.
-Experiments and Discussion
-In our experiments, we use a toy denoising example to validate our theorems and demonstrate the proposed training scheme using seismic deconvolution and magnetic resonance imaging (MRI) reconstruction.
-For each task, we train the same LU architecture using different training schemes.
-We then evaluate the models using in-distribution testing data, adversarial attacks with various strengths, and task-dependent OOD data.
-Detailed training procedures are provided in the Appendix.
-Toy problem: 2-dimensional Point Denoising
-To validate the Theorems, we compare the robustness and generalization using regular MSE, AT and SGD-jittering in a point denoising problem.
-Assume ${\bm{x}}=(0,0)^{\top}$, we generate noisy ${\bm{y}}$ by adding random zero-mean Gaussian noise with a variance of 0.01.
-We use 200 samples for training and 50 for evaluation.
-Figure 1 shows the noisy points ${\bm{y}}$ and the denoised results $\hat{{\bm{x}}}$ (blue) across four evaluation modes, with the ground-truth in red.
-In-distribution data are sampled from the same distribution as the training data, centered at $(0,0)^{\top}$.
-The worst-case attack finds a vector within a 0.01 $\ell_{2}$ distance around ${\bm{y}}$ that maximizes the squared error, while the average-case attack uniformly samples attack vectors within the same ball and adds to ${\bm{y}}$.
-SGD jittering and AT are more robust than MSE training in both metrics.
-AT aims to promote worst-case robustness, which implies average-case robustness for attacks within the same $\ell_{2}$ ball around ${\bm{y}}$, though its average-case performance may not be optimal.
-We observe that SGD jittering has even tighter clusters in its reconstructions under average-case attacks than AT.
-Finally, we evaluate generalization accuracy by reconstructing the slightly biased ground truth $(0.005,0)^{\top}$ from its corresponding measurements.
-With the introduction of bias, SGD jittering achieves a more accurate reconstruction compared to the new ground truth.
-Notice that MSE training underperforms on in-distribution data compared to SGD jittering, likely because the noise injection in SGD jittering helps the model escape from local minima in training.
-Figure 3: Seismic deconvolution with an additional layer.
-Figure 4: MRI evaluation (a) under adversarial attack obtained from projected gradient descent with norm no more than $\epsilon=1$.
-Each row represents an MRI reconstruction of a sample from fastMRI (Zbontar et al.
-2018a).
-(b) MRI evaluation using OOD tumor knee data from (Bickle and Jin 2021).
-Seismic Deconvolution
-Seismic deconvolution is a crucial problem in geophysics.
-It aims to reverse the convolution effects in the received signal, which occurs when artificial source waves travel through the Earth’s layers.
-The goal is to extract sparse reflectivity series from recorded seismic data.
-The training data is generated following the same procedure as in (Iqbal et al.
-2019; Guan et al.
-2024).
-To demonstrate the generalization accuracy, we generate OOD testing data by adding an arbitrary horizontal layer with magnitudes of 0.1 and 0.3 to the ground truth, reflected accordingly in the measurements.
-Table 1 numerically compares the average testing peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) of the reconstructions.
-Notice that for OOD data, the numerical values are very similar, as the additional layer has a small magnitude relative to the overall measurement.
-However, the difference can be clearly observed in the figures.
-Figure 2 shows two deconvolution examples for each trained model under adversarial attacks.
-MSE training produces artifacts, while both AT and SGD jittering yield results closer to the ground truth.
-Notably, in the top example, AT fails to recover the small horizontal layer in the center of the boxed area, but SGD jittering successfully captures it.
-Figure 3 demonstrates the generalization ability with two different magnitude levels.
-For the OOD example with the larger magnitude, both SGD jittering and MSE training recover a cleaner layer compared to AT.
-More notably, with the smaller magnitude, AT perceives the line as an unwanted perturbation and barely recovers it, whereas the result from SGD jittering is more distinctly visible.
-PSNR / SSIM Test data Adv. Attack OOD small OOD large
-MSE training 34.64 / 0.921 28.85 / 0.829 34.57 / 0.918 34.57 / 0.917
-AT 33.50 / 0.903 30.27 / 0.849 33.45 / 0.902 33.43 / 0.902
-Input Jittering 32.92 / 0.882 30.10 / 0.832 32.91 / 0.882 32.90 / 0.882
-SGD jittering 35.10 / 0.928 29.89 / 0.842 34.93 / 0.927 34.93 / 0.927
-Table 1: Seismic deconvolution evaluation for in-distribution data (column 2), under adversarial attack with $\epsilon=1$ (column 3), and for OOD data with small and large magnitudes (columns 4 and 5).
-The best and second-best performances are in bold and underlined respectively.
-Accelerated MRI Reconstruction
-Accelerated MRI aims to recover human-interpreted body structures from partial k-space measurements.
-We train models with single-coil knee MRI from the fastMRI dataset (Zbontar et al.
-2018b) with 4×4$\times$4 × acceleration, or $1/4$ of the measurements in k-space is used for reconstruction.
-To assess generalization, we use a different knee dataset from Bickle and Jin (2021), which includes giant tumor cells not present in the training dataset.
-Table 2 numerically compares the average testing PSNR and SSIM of the reconstructions for in-distribution fastMRI data, under adversarial attack with $\epsilon=1$ and OOD tumor cell knee data.
-Our findings indicate that MSE training and SGD jittering exhibit comparable performance on in-distribution data, but SGD jittering demonstrates greater robustness and superior performance on OOD data.
-Adversarial training (AT) yields the most robust model under attack, though it comes with a significant tradeoff in performance for both in-distribution and OOD data.
-Additionally, as shown in Figure 4, reconstructions using AT appear blurrier than others, a similar observation noted in black-box IP solvers by Krainovic, Soltanolkotabi, and Heckel (2023).
-PSNR / SSIM fastMRI Adv. Attack Tumor Cell
-MSE training 28.21 / 0.603 25.68 / 0.382 29.92 / 0.779
-AT 27.68 / 0.564 27.17 / 0.549 27.74 / 0.597
-Input Jittering 28.18 / 0.595 25.05 / 0.420 29.97 / 0.740
-SGD jittering 28.22 / 0.607 26.77 / 0.552 30.36 / 0.788
-Table 2: MRI evaluation for in-distribution data (column 2), adversarial attack (column 3), and OOD data (column 4). The best and second best performances are in bold and underlined respectively.
-Conclusion
-Robustness and generalization accuracy are both crucial in solving IPs, yet current analyses often focus on simple black-box solvers that enhance robustness at the cost of accuracy.
-While empirical studies have examined the robustness of MBAs, there is a lack of theoretical analysis.
-In this work, we introduce a novel and easily implemented SGD jittering training scheme to address the tradeoff between robustness and accuracy in solving IPs.
-Our method proves effective for high-quality model-based IP solvers.
-Through mathematical analysis, we demonstrate that this approach implicitly regularizes the gradient and Hessian of the neural network with respect to the input, resulting in improved generalization and enhanced robustness compared to traditional MSE training.
-Experimental results show that while MSE training is vulnerable to adversarial attacks and AT recovers smooth estimates, SGD jittering consistently produces robust, high-quality outcomes.
-Appendix A Appendix
-Appendix B Notations
-We clarify the notations used in the paper.
-  • $H_{\theta}$: learned inverse mapping parameterized by $\theta$, i.e., black-box neural networks, MBAs.
-For MBAs in particular, $H_{\theta}=\operatorname*{arg\,min}_{x}~{}\frac{1}{2}\|{\bm{y}}-\bm{A}{\bm{x}}\|^{2}_{2}+r_{\theta}({\bm{x}})$.
-  • $r$: regularization function from predefined functions, i.e., $\ell_{1}$ and $\ell_{2}$ norm.
-  • $r_{\theta}$: regularization function whose gradient is learned from a neural network $f_{\theta}$.
-  • $f_{\theta}$: learned gradient update of $r$, $f_{\theta}({\bm{x}})=\nabla_{\bm{x}}r_{\theta}({\bm{x}})$.
-  • $F(x)=\frac{1}{2}\|{\bm{y}}-\bm{A}{\bm{x}}\|^{2}_{2}+r_{\theta}({\bm{x}})$, denote the value of the lower-level objective function.
-Appendix C Proof of Lemma 1
-Lemma 4.
-(Restate Lemma 1.
-Convergence of MBAs-SGD for IPs) Let $F({\bm{x}})=\frac{1}{2}||{\bm{y}}-\bm{A}{\bm{x}}||_{2}^{2}+r({\bm{x}})$ denote the lower-level objective function.
-Assume $f_{\theta}({\bm{x}})$ is $L$-Lipschitz continuous,
-$||f_{\theta}({\bm{x}}+\Delta)-f_{\theta}({\bm{x}})||_{2}\leq L||\Delta||_{2},\quad\forall{\bm{x}},\Delta$
-With assumption 1, and step-sizes $\eta_{k}<2/L$, SGD in minimizing $F({\bm{x}})$ converges to a stationary point.
-Proof.
-The lemma can be easily derived from the Theorem in (Ghadimi and Lan 2013), which assumes $\nabla_{\bm{x}}F({\bm{x}})$ is $L$-Lipschitz continuous.
-$||\nabla_{\bm{x}}F({\bm{x}}+\Delta)-\nabla_{\bm{x}}F({\bm{x}})||_{2}$
-$=$ $||\bm{A}^{\top}(\bm{A}({\bm{x}}+\Delta)-{\bm{y}})+f_{\theta}({\bm{x}}+\Delta)-\bm{A}^{\top}(\bm{A}{\bm{x}}-{\bm{y}})-f_{\theta}({\bm{x}})||_{2}$
-$=$ $||\bm{A}^{\top}\bm{A}\Delta+f_{\theta}({\bm{x}}+\Delta)-f_{\theta}({\bm{x}})||_{2}$
-$\leq$ $||\bm{A}^{\top}\bm{A}||\,||\Delta||_{2}+||f_{\theta}({\bm{x}}+\Delta)-f_{\theta}({\bm{x}})||_{2}$
-$\leq$ $\max_{i=1,..,n}\lambda_{i}||\Delta||_{2}+||f_{\theta}({\bm{x}}+\Delta)-f_{\theta}({\bm{x}})||_{2}$
-where $\lambda_{i}$ be the eigenvalues of $\bm{A}^{\top}\bm{A}$, and let $\mu=\max_{i=1,..,n}\lambda_{i}$.
-Since $\nabla F$ is $L$-Lipshitz, we show that for MBAs-SGD in solving denoising problems, it is equivalent to $(L-\mu)$-Lipschitz continuous of $f_{\theta}$.
-With the same assumptions on step sizes and the SGD jittering noise, we finish the proof.
-∎
-Appendix D Proof of Theorem 3 - robustness
-Theorem 5.
-(Restate Theorem 3) Let ${\bm{e}}$ be the perturbation vector and let $\mathcal{P}_{e}$ be iid zero-mean Gaussian distribution such that $\mathbb{E}||{\bm{e}}||_{2}^{2}=\sigma_{\bm{e}}^{2}$.
+Appendix D Proof of Robustness (Theorem 7.5)
+Theorem D.1.
+(Restate Theorem 7.5) Let ${\bm{e}}$ be the perturbation vector and let $\mathcal{P}_{e}$ be iid zero-mean Gaussian distribution such that $\mathbb{E}||{\bm{e}}||_{2}^{2}=\sigma_{\bm{e}}^{2}$.
 The average-case robustness risk for denoising problem is as follows,
-$R_{e}(\theta)=\mathbb{E}_{({\bm{x}},{\bm{y}})\in\mathcal{D},{\bm{e}}}~{}\left[||{\bm{x}}-H_{\theta}({\bm{y}}+{\bm{e}})||^{2}_{2}\right].$ (10)
+$R_{e}(\theta)=\mathbb{E}_{({\bm{x}},{\bm{y}})\sim\mathcal{D},{\bm{e}}\sim\mathcal{P}_{e}}~{}\left[||{\bm{x}}-H_{\theta}({\bm{y}}+{\bm{e}})||^{2}_{2}\right].$ (14)
 For the denoising problem, SGD jittering is more robust than MSE training against bounded-variance perturbations around the measurement, or
@@
 $R_{e}(\theta)=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}-{\bm{e}}-\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))||^{2}_{2}$
-$=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))||^{2}_{2}$
-$\quad+\mathbb{E}||{\bm{e}}||_{2}^{2}+2\mathbb{E}\left[<{\bm{e}},\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))>\right]$
-$=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))||^{2}_{2}$
-$\quad+\mathbb{E}||{\bm{e}}||_{2}^{2}+2\mathbb{E}\left[<{\bm{e}},-\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}-f_{\theta}({\bm{x}}_{i}^{at})>\right]$
-$=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))||^{2}_{2}+\sigma_{\bm{e}}^{2}$
-$\quad+2\mathbb{E}\left[\langle{\bm{e}},-\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}-f_{\theta}({\bm{x}}_{i}^{at})\rangle\right]$
+$=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))||^{2}_{2}+\mathbb{E}||{\bm{e}}||_{2}^{2}+2\mathbb{E}\left[\langle{\bm{e}},\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))\rangle\right]$
+$=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))||^{2}_{2}+\sigma_{\bm{e}}^{2}+2\mathbb{E}\left[\langle{\bm{e}},\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))\rangle\right]$
 The first term is the mean-squared testing loss for in-distribution data.
@@
 The last term also appears in SGD jittering risk, when the above condition for jittering level and robustness level are satisfied, ${\bm{e}}$ and $\sum_{i=0}^{K}\eta(1-\eta)^{K-i}{\bm{w}}_{i}$ are from the same distribution.
-Therefore, for models that perform equally well in mean-squared error for in-distribution testing data, minimizing SGD jittering risk minimizes overall robustness.
+Therefore, for models that perform equally well in MSE for in-distribution testing data, minimizing SGD jittering risk minimizes overall robustness.
 ∎
-Appendix E Training Details
+Appendix E Extra Stochastic Proximal Gradient Descent Experiments
+Similar to gradient descent variants of loop unrolling architectures, for proximal gradient descent (PGD) variants, we train a PGD LU with regular MSE loss.
+Then finding the worst-case attack to ${\bm{y}}$ over all PGD updates.
+E.1 Seismic Deconvolution
+In the same seismic deconvolution problem setup discussed in the main manuscript, but now solved using proximal gradient variants, Figure 7(a) shows the performance of the trained models under an adversarial attack with $\epsilon=1$.
+Figure 7(b) focuses on reconstructing the additional layer highlighted in red boxes.
+Table 3 provided numerical evaluations.
+The proposed SPGD jittering demonstrates greater robustness compared to MSE training while better preserving detailed information than adversarial training (AT).
+Figure 7: Visualization of seismic deconvolution using (stochastic) proximal gradient descent variants within loop unrolling architectures.
+Figure (a) illustrates the reconstruction performance under adversarial attack with $\epsilon=1$.
+Figure (b) evaluates the model’s performance on out-of-distribution (OOD) seismic data, where the ground-truth reflectivity includes an additional small horizontal layer, highlighted in the red boxes.
+PSNR / SSIM in-distribution Adv. Attack OOD
+MSE training 34.91 / 0.923 28.78 / 0.842 34.810 / 0.920
+AT 32.99 / 0.899 31.44 / 0.875 32.966 / 0.899
+Input Jittering 33.02 / 0.899 30.56 / 0.861 32.910 / 0.895
+SPGD Jittering 35.05 / 0.931 29.93 / 0.870 34.952 / 0.929
+Table 3: PGD variants of LU are evaluated on seismic deconvolution across three scenarios: in-distribution data (column 2), adversarial attacks (column 3), and out-of-distribution (OOD) data (column 4).
+The best performance in each case is highlighted in bold, while the second-best performance is underlined.
+E.2 Accelerated MRI Reconstruction
+We complete the missing numerical results discussed in the main manuscript for 4×4$\times$4 × accelerated MRI reconstruction using SPGD jittering in Table 4.
+PSNR / SSIM fastMRI Adv. Attack Tumor Cell
+MSE training 27.93 / 0.582 22.82 / 0.179 29.39 / 0.771
+AT 27.20 / 0.533 26.78 / 0.518 27.13 / 0.560
+Input Jittering 24.88 / 0.299 24.44 / 0.297 25.85 / 0.478
+SPGD Jittering 28.15 / 0.597 27.56 / 0.577 29.762 / 0.777
+Table 4: PGD variants of LU are evaluated on MRI data across three scenarios: in-distribution data (column 2), adversarial attacks (column 3), and out-of-distribution (OOD) data (column 4).
+The best performance in each case is highlighted in bold, while the second-best performance is underlined.
+Appendix F Compared to Other Baselines for MRI reconstruction
+We evaluate our method against a widely used robustness technique Lipschitz regularization using spectral normalization (SN).
+SN constrains the Lipschitz constant of the network by bounding the spectral norm (i.e., largest singular value) of each weight matrix, which helps stabilize training and improve robustness.
+As shown in the second last row in Table 5 for 4xMRI reconstruction.
+SN promotes adv.
+robustness, but at the cost of reduced accuracy to some extend.
+This is likely due to its restrictive nature of SN limiting the model’s expressive power.
+Additionally, we also compares the performance to a diffusion models (DMs), as they also demonstrates impressive results in image generation.
+Since our work proposes a general framework for IPs, we chose to compare against the standard DDPM rather than specialized task-specific DMs, consistent with prior work (Güngör et al., 2023).
+To ensure a fair comparison under similar computational constraints, we adapt the denoising U-Net to fit within the same GPU memory as other methods.
+Table 5 compares DDPM with other methods.
+DDPM achieves comparable in-distribution (ID) performance to MBAs trained with both MSE loss and the proposed SGD jittering, and shows stronger robustness than standard MSE-trained MBAs.
+However, it underperforms in OOD generalization.
+While DDPM serves as a strong baseline for robustness and ID accuracy, SGD jittering achieves better generalization under distribution shifts.
+We will add the comparison to DM in the main manuscript as an interesting baseline.
+It is also worth noting that DDPM requires  10× more parameters and is significantly more data-intensive, whereas MBAs are more data-efficient (Monga et al., 2021) due to their optimization-inspired iterative structure.
+We also refer to prior work (Güngör et al., 2023), which compares DDPM, DiffRecon, AdaDiff to MSE-trained MBAs for MRI reconstruction.
+Their results show that while DM can generalize well in some cases, MBA methods consistently perform better on ID data.
+Results in (Güngör et al., 2023) shows that MSE-trained MBA is a strong baseline, and our proposed SGD jittering further improves their robustness and generalization.
+Model Training Scheme/Design fastMRI Adv. Attack Tumor Cell
+GD LU MSE training 28.21 / 0.603 25.68 / 0.382 29.92 / 0.779
+AT 27.68 / 0.564 27.17 / 0.549 27.74 / 0.597
+Input Jittering 28.18 / 0.595 25.05 / 0.420 29.97 / 0.740
+SGD jittering (Ours) 28.22 / 0.607 26.77 / 0.552 30.36 / 0.788
+Lipschitz regularization (SN) 27.71 / 0.576 27.09 / 0.542 27.92 / 0.594
+DDPM Default training 28.17 / 0.611 27.13 / 0.536 29.72 / 0.782
+Table 5: MRI reconstruction evaluation compared to the Lipschitz regularization method (last row). The best and second best performances are in bold and underlined respectively.
+Appendix G Training Details
 We use 10-iteration gradient descent loop unrolling architectures for all tasks.
@@
 In input jittering training, the three methods in order of 2D denoising, seismic deconvolution and MRI reconstruction use the noise variance of 0.01, 0.05 and 0.05.
-In SGD jittering training, we choose the DGD jittering variance for each task with 0.01, 0.1 and 0.01 respectively.
-Appendix F Training Speed
-We also compare the training speed among different training schemes.
+In SGD jittering training, we choose the SGD jittering variance for each task with 0.01, 0.1 and 0.01 respectively.
+Appendix H Training Speed
+We also compare the training speed among different training schemes for GD and PGD LU architectures.
 All methods are trained on the same device.
@@
 AT is trained by projected gradient descent.
-For the 2D denoising problem, $\epsilon=0.01$ with 50 PGD iterations and a stepsize of 0.05.
-For the seismic deconvolution and MRI reconstruction problem, $\epsilon=1$, with 20 PGD iterations and a stepsize of 0.1.
+For the 2D denoising problem, $\epsilon=0.01$ with 50 projected GD iterations and a stepsize of 0.05.
+For the seismic deconvolution and MRI reconstruction problem, $\epsilon=1$, with 20 projected GD iterations and a stepsize of 0.1.
+AT is significantly slower than other methods.
+Notice that the jittering in SGD and SPGD does introduce very minimal time overhead compared to standard MSE training, even when accounting for the additional noise sampling steps.
 2D denoising Seis. Deconv. MRI
-MSE training 5360.84 it/s 23.69 it/s 5.23 it/s
+GD LU MSE training 5360.84 it/s 23.69 it/s 5.23 it/s
 AT 562.75 it/s 1.53 it/s 0.25 it/s
@@
 SGD jittering 4337.84 it/s 22.72 it/s 5.27 it/s
+PGD LU MSE training 5136.34 it/s 16.78 it/s 6.23 it/s
+AT 486.68 it/s 1.12 it/s 0.17 it/s
+Input Jittering 4965.25 it/s 16.44 it/s 6.17 it/s
+SPGD jittering 4669.76 it/s 14.52 it/s 6.07 it/s
 Table 6: Training speed measured in items per second are recorded.
-AT is significantly slower than other methods. Notice that SGD jittering is very time efficient, even with the additional noise sampling steps.